### v1

# UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models

Abstract
Large language models require iterative updates to address challenges such as knowledge conflicts and outdated information (e.g., incorrect, private, or illegal contents). Machine unlearning provides a systematic methodology for targeted knowledge removal from trained models, enabling elimination of sensitive information influences. However, mainstream fine-tuning-based unlearning methods often fail to balance unlearning efficacy and model ability, frequently resulting in catastrophic model collapse under extensive knowledge removal. Meanwhile, in-context unlearning—which relies solely on contextual prompting without modifying the model’s intrinsic mechanisms—suffers from limited generalizability and struggles to achieve true unlearning. In this work, we introduce UniErase, a novel unlearning paradigm that employs learnable parametric suffix (unlearning token) to steer language models toward targeted forgetting behaviors. UniErase operates through two key phases: (I) an optimization stage that binds desired unlearning outputs to the model’s autoregressive probability distribution via token optimization, followed by (II) a lightweight model editing phase that activates the learned token to probabilistically induce specified forgetting objective. Serving as a new research direction for token learning to induce unlearning target, UniErase achieves state-of-the-art (SOTA) performance across batch, sequential, and precise unlearning under fictitious and real-world knowledge settings. Remarkably, in terms of TOFU benchmark, UniErase, modifying only around 3.66% of the LLM parameters, outperforms previous forgetting SOTA baseline by $\sim$ 4.01$\times$ for model ability with even better unlearning efficacy. Similarly, UniErase, maintaining more ability, also surpasses previous retaining SOTA by 35.96% for unlearning efficacy, showing dual top-tier performances in current unlearing domain. We release our code at  https://github.com/Ymm-cll/UniErase.
1 Introduction
While the large language models (LLMs) community xu2024survey ; kumar2024large ; touvron2023llama ; guo2025deepseek ; chang2024survey ; wang2025comprehensive has made significant advances in enhancing the reasoning and memorization capabilities through pretraining ma2024megalodon ; mckinzie2024mm1 ; tirumala2023d4 and post-training kumar2025llm ; gan2024towards ; huang2024billm ; tie2025survey techniques, the complementary challenges of enabling precise and efficient unlearning remain crucial to address privacy, ethical and security concerns lu2024eraser ; zhang2024safe ; xu2024machine ; si2023knowledge . The heterogeneous nature of training data sources and massive parameters of LLMs introduce significant difficulties in controlling the generation of sensitive, private, or illegal contents yao2024survey ; shi2024large . These privacy-related risks and potential misuse cases have made all individual entities very cautious when deploying LLMs – let alone large enterprises and governmental organizations he2024emerged ; wang2025comprehensive ; yu2025survey .
Striking an optimal balance between model ability and unlearning efficacy remains fundamentally challenging geng2025comprehensive ; qu2024frontier . The core objective of unlearning is to enable LLMs trained on trillion-token corpora to completely forget specific data subsets while preserving retaining data and their general capabilities si2023knowledge ; maini2024tofu . Instead of prohibitively expensive retraining, fine-tuning (FT) via carefully designed loss functions is the current mainstream methods, which can be broadly categorized into two types: (I) Targeted unlearning involves deliberately erasing certain data or learned associations in LLMs to modify their outputs in a controlled and specified manner zhang2024negative ; wei2021finetuned ; rafailov2023direct ; maini2024tofu ; yuan2024closer . (II) Untargeted unlearning shifts the responses away from the original outputs but without specifying a particular direction maini2024tofu ; zhang2024negative ; yuan2024closer ; wang2025uipe ; wang2025rethinking . Typically, this line employs techniques such as gradient ascent maini2024tofu and negative preference optimization zhang2024negative to steer the model away from knowledge in the forget set. Notably, in-context unlearning (ICU) is a complementary paradigm that approximates unlearning through prompt engineering alone, without modifying the model’s parameters pawelczyk2023context ; takashiro2024answer ; bhaila2024soft .
Figure 1: Our UniErase achieves comprehensive and balanced unlearning (Left) while maintains consistently high LLM capability (Middle), delivering rapid processing and high precision (Right).
However, the observed empirical evidence gaps persist across the above methods, indicating non-trivial challenges in current unlearning frameworks geng2025comprehensive ; maini2024tofu ; yuan2024closer . Specifically, ICU-based methods suffer from poor generalizability, robustness, and adaptability pawelczyk2023context —for example, they can only handle unlearning in classification tasks and fail to achieve true forgetting. Meanwhile, FT-based approaches struggle to strike a balance between unlearning efficacy and model ability maini2024tofu ; qu2024frontier , with other problems like over-unlearning and unpredictable behaviors yuan2024closer . Empirical analysis reveals a critical limitation, for batch unlearning, the existing best method in forgetting loses 80.28% of the LLM’s general ability, while the retaining best only forgets around half of the required data.
Generally, when predicting the next token’s probability distribution, LLMs are heavily influenced by the preceding token sequences. Specifically, in the safety domain, the affirmative or negative tone of a model’s initial response can significantly alter the success rate of replying to harmful queries zou2023universal ; qi2024safety ; dong2024attacks . Then, for the unlearning scenario, a natural but non-simple question arises: Can an unlearning token, denoted as [UNL], redirect the LLM’s “cognition” of the forgetting set $\mathcal{D}_{f}$ toward an unlearning space while preserving the response direction for the retaining set $\mathcal{D}_{r}$?
To this end, we propose UniErase, a universal unlearning method that balances unlearning efficacy and model ability. Technically, UniErase mainly consists of two parts: unlearning token and unlearning editing. Inspired by meta-tokens li2021prefix ; lester2021power , UniErase first learns an unlearning token (denoted as [UNL]) on the forgetting set $\mathcal{D}_{f}$ and ensures other LLM parameters remaining unaffected by freezing them. To learn [UNL], we only introduce two new vectors in the embedding and unembedding matrices (projecting hidden states into next-token probability distribution). Through optimization on these two vectors, the LLM is guided to output “I don’t know”-like responses when [UNL] following input knowledge (in $\mathcal{D}_{f}$) queries. Notably, without [UNL], the LLM still responds $\mathcal{D}_{f}$ and the retaining set $\mathcal{D}_{r}$ normally. Thus, to build the connection between [UNL] and $\mathcal{D}_{f}$, we further propose unlearning editing, which modifies a minimal amount of model parameters to steer $\mathcal{D}_{f}$ into the “unlearning space” via [UNL] and preserves knowledge in $\mathcal{D}_{r}$. Specifically, by adjusting a few parameters in specific Transformer layers, for the knowledge queries in $\mathcal{D}_{f}$, we alter the first token of their responses to [UNL], leveraging its properties to ensure the model forgets these knowledge. In a parallel vein, we maintain the knowledge associations in $\mathcal{D}_{r}$ to keep the LLM’s general ability.
We conduct extensive experiments to validate the effectiveness of our UniErase. In fact, we consider fictitious and real-world knowledge in batch, sequential and precise unlearning scenarios for the Llama series of LLMs. Evaluating via multi-dimensional metrics on TOFU benchmark, UniErase significantly outperforms the previous SOTA forgetting baseline, achieving a 4.01$\sim$ 4.01$\times$ 4.01 × improvement in model ability while demonstrating better unlearning. Additionally, compared with prior retaining SOTA, UniErase maintains superior model ability and is 35.96% higher in forgetting efficacy.
In summary, our contributions can be listed as follows:
  • Brand-new Paradigms. Building upon UniErase, we propose a novel paradigm that achieves efficient unlearning by directly modifying model parameters instead of multi-round fine-tuning, significantly expanding the scope of future research in the current unlearning technology stack.
  • Balanced Trade-off. Compared to previous SOTA methods, UniErase achieves more thorough unlearning and better general ability preserving, boosting the practical usability of LLM unlearning.
  • Generalized Scenarios. UniErase achieves the best performances across batch, sequential and precise unlearning for fictitious and real-world knowledge, covering nearly all unlearning cases.
2 Related Works
Machine Unlearning. The concept of machine unlearning from traditional models bourtoule2021machine ; chen2022graph ; nguyen2022survey is emerging as a rising approach for LLM alignment research liu2025rethinking ; thaker2024position . Its primary goal is to enable LLMs to forget a subset $\mathcal{D}_{f}$ of training data $\mathcal{D}$ (e.g., privacy or harmful knowledge) and maintain the abilities on a retain set $\mathcal{D}_{r}\subset\mathcal{D}$, without the high cost of retraining geng2025comprehensive . Current mainstream approaches relying on fine-tuning and designing various loss functions for different objectives (forget $\mathcal{D}_{f}$ and retaining $\mathcal{D}_{r}$). Specifically, GD liu2022continual reduces the probability of generating responses to $\mathcal{D}_{f}$ by ascending gradients, and introduces another loss to constrain the deviation between the unlearned model and the original one liu2022continual . Meanwhile, NPO zhang2024negative , based on preference optimization rafailov2023direct , achieves unlearning by solely using $\mathcal{D}_{f}$ as negative preference, ignoring the positive term. Other works, such as RMU huu2024effects and LUNAR shen2025lunar , employ steering-vector-like approaches cao2024personalized to forcibly modify hidden states (but still via fine-tuning) and redirect $\mathcal{D}_{f}$ toward inability space. Additionally, SPULbhaila2024soft makes preliminary attempts in unlearning by adding soft prompts li2021prefix ; lester2021power during inference to manipulate model responses, but without modifying parameters to achieve “essential forgetting”.
Model Editing. LLMs may contain outdated or even incorrect information huang2025survey ; tonmoy2024comprehensive . However, like unlearning, retraining to remove is costly, while fine-tuning overfits with limited updates. Thus, the light and precise model editing wang2024knowledge ; he2024llms is proposed for truthfulness huang2024can , and safety li2024precision ; chen2024can ; li2024badedit . Specifically, interpretability study treats LLMs’ two-layer Multilayer Perceptrons (MLP) as lookup tables for storing knowledge geva2020transformer ; zhao2021non . So Early methods like ROME meng2022locating and MEMIT meng2022mass introduce the locate-then-edit paradigm by solving closed-form optimization on the MLP’s down-projection matrix. AlphaEdit fang2024alphaedit further efficiently preserves other knowledge by projecting updates into the null space of the original knowledge representation. AnyEdit jiang2025anyedit decomposes edits into sub-operations for more complex knowledge. However, editing itself cannot accomplish unlearning directly: it requires knowledge represented in (subject, relation, object) triples and updates the object to a new value, but there exists no object token that corresponds to the abstract concept of unlearning. Our UniErase introduces unlearning-specific editing that integrates with existing methods.
3 Preliminaries
Notations. We consider an LLM as a probabilistic language model with parameters $\theta$, denoted as $\pi_{\theta}$. The target set for forgetting and the set of retained knowledge are represented as $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, where typical elements of both are question and answer pairs in the form of $d=(q,a)$. In addition, we denote the set of real numbers as $\mathbb{R}$, and the set of real number tensors with dimensions $(d_{1},...,d_{n})$ as $\mathbb{R}^{d_{1}\times...\times d_{n}}$. The symbol $\oplus$ represents the concatenation operation of token sequences.
Unlearning Target. For an LLM $\pi_{\theta}$ trained with dataset $\mathcal{D}$, unlearning process aims to make the model forget the contents in $\mathcal{D}_{f}\subset\mathcal{D}$, such that it behaves as if it were trained solely on $\mathcal{D}\setminus\mathcal{D}_{f}$. In a parallel vein, unlearning must preserve the model’s knowledge on other datasets and its general capabilities. Similar to the trade-off between safety and helpfulness in LLM optimization varshney2023art , unlearning involves a balance between unlearning efficacy and model ability, formulated as:
$\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}\big{[}\sum_{d\in\mathcal{D}_{f}}\mathrm{Forget}(d;\pi_{\theta})+\sum_{d\in\mathcal{D}_{r}}\mathrm{Ability}(d;\pi_{\theta})\big{]},$ (1)
where “$\mathrm{Forget}$” and “$\mathrm{Ability}$” are the standards or metrics for unlearning efficacy and model ability.
Mainstream Unlearning Strategies. To achieve the goal in Eq 1, mainstream unlearning methods design distinct forgetting/retaining losses $l_{f}/l_{r}$ on $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, sometimes using the original model $\pi_{\theta}^{\text{ref}}$ as a reference. We unify their loss designs as follows, with $\beta$ and $\gamma$ as the weights:
$\arg\min_{\pi_{\theta}}=\beta\underbrace{\mathbb{E}_{(q,a)\sim\mathcal{D}_{f}}\big{[}l_{f}(q\mid a;\pi_{\theta},\pi_{\theta}^{\text{ref}})\big{]}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{forgetting term}}}+\gamma\underbrace{\mathbb{E}_{(q,a)\sim\mathcal{D}_{r}}\big{[}l_{r}(q\mid a;\pi_{\theta},\pi_{\theta}^{\text{ref}})\big{]}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{retaining term}}}.$ (2)
In Eq 2, the forgetting term is designed to make the model forget the content on $\mathcal{D}_{f}$, while the retaining term aims to preserve the knowledge on $\mathcal{D}_{r}$. However, current methods typically select the neighboring knowledge of $\mathcal{D}_{f}$ as $\mathcal{D}_{r}$, and can not encompass diverse general knowledge due to high time cost. In Appendix B, we introduce the specific forms of various $l_{f}$ and $l_{r}$ in detail.
4 Methodology
In this section, we initially employ mathematical logic to expound upon the fundamental principles of UniErase ($\triangleright$ Sec 4.1), as demonstrated in Figure 2. Then, we elaborate on the methodology of training an unlearning token via incorporating a minimal number of parameters ($\triangleright$ Sec 4.2). Additionally, we discuss the process of precisely modifying parameters for the forgetting targets ($\triangleright$ Sec 4.3), aiming at achieving an optimal equilibrium between high unlearning efficacy and low ability loss ($\triangleright$ Sec 4.4).
4.1 Unlearning Logical Chain
Given an LLM $\pi_{\theta}$, for an input token sequence $q=[x_{1}x_{2}...x_{n}]$, we assume the output token sequence is $a=[y_{1}y_{2}...y_{m}]$. Then we abstract this generation process as mathematical logic derivations:
$x_{1}x_{2}...x_{n}\xRightarrow{\pi_{\theta}}y_{1}\xRightarrow{\pi_{\theta}}y_{2}\xRightarrow{\pi_{\theta}}...\xRightarrow{\pi_{\theta}}y_{m},$ (3)
where each $\xRightarrow{\pi_{\theta}}$ represents generating the next token based on all previously generated tokens.
Proposition 1. The Unlearning Token is a novel token trained on $\mathcal{D}_{f}$, designed to constrain $\pi_{\theta}$ such that its subsequent token generation redirects targeted concept forgetting or induced unawareness. Specifically, for $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$, the unlearning token [UNL] is defined as follows:
$x_{1}x_{2}...x_{n}\oplus\mathrm{[UNL]}\xRightarrow{\pi_{\theta}}y_{\text{idk}},\quad y_{\text{idk}}\in\mathcal{D}_{\text{idk}},$ (4)
where $\mathcal{D}_{\text{idk}}$ is the set of token sequences that express forgetting or ignorance (e.g., “I do not know”).
Building upon Proposition 1, we have learned the [UNL] token. However, when only $q$ is provided as input, the model still generates original normal response rather than $y_{\text{idk}}$. Therefore, we need to modify model parameters so that for any $q\in\mathcal{D}_{f}$, the first token of the generated response is [UNL], thereby internalizing “forgetting $q$ with [UNL]” as part of the model’s knowledge or “cognition”. Model editing, as an evolving approach to edit model knowledge without ability degradation fang2024alphaedit ; meng2022locating , basically meets our unlearning requirements if adapted properly. To this end, we propose:
Proposition 2. Unlearning Editing utilizes model editing to modify only a small set of parameters ($|\Delta\theta|\ll|\theta|$), enabling the model to forget specified intrinsic knowledge. For $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$ with the obtained [UNL], the parameter $\theta$ is updated with $\Delta\theta$, formally defined as:
$\theta\leftarrow\theta+\Delta\theta,\quad|\Delta\theta|\ll|\theta|,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}.$ (5)
Derivation: Eq 4 ensures that [UNL] can direct the input to the unlearning space (generating responses in $\mathcal{D}_{\text{idk}}$) when used as a suffix, while Eq 5 changes the next token of the normal response of the input sequence to [UNL]. Based on these two propositions, we can derive the following Unlearning Logical Chain to adjust model parameters for achieving our UniErase unlearning:
$\theta\leftarrow\theta+\Delta\theta,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\xRightarrow{\mathrm{\pi_{{\theta}}}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}.$ (6)
This chain enables us to redirect $\mathcal{D}_{f}$ into the unlearning space and leverage model editing to preserve general capabilities. However, the following challenges arise for implementation:
Challenge I: How to teach LLMs to learn and generate [UNL] without overfitting? Since the off-the-shelf soft prompt li2021prefix ; lester2021power ; bhaila2024soft optimization only considers meta tokens as input prefix or suffix without truly introducing a new token. To represent the unlearning space, we need to teach the model the meaning of the unlearning token [UNL]. Besides, editing in Eq 5 requires LLM’s ability to generate the target token (as one step involves maximizing its generation probability). Therefore, we should optimize the model to generate [UNL]. However, the samll $|\mathcal{D}_{f}|$ may lead to overfitting, where the LLM incorrectly holds $q\xRightarrow{\pi_{\theta}}\mathrm{[UNL]}$ for $(q,a)\notin\mathcal{D}_{f}$ (even for $(q,a)\in\mathcal{D}_{r}$) before editing.
Challenge II: How to enhance the robustness of [UNL]? The above derivation is based on model’s continuing generation, but aligned LLMs in practical applications typically append chat templates to input (e.g., user/assistant role, demonstration in Appendix C), which may cause [UNL] fail to induce $\mathcal{D}_{\text{idk}}$. On the other hand, in Eq 5, even if the scale of parameter change is small ($|\Delta\theta|\ll|\theta|$), $\Delta\theta$ may turn Eq 4 not true for the updated model. Thus, for the derivation in Eq 6, we need to ensure that [UNL] remains effective after parameter modifications, successfully serving as a token to induce $y_{\text{idk}}$.
Figure 2: Overview framework of UniErase. We demonstrate current unlearning methods and our UniErase, consisting of unlearning logical chain, unlearning token and unlearning editing.
4.2 Generating Unlearning Token (Proposition 1)
In this section, with Figure 2, we first introduce how to learn the unlearning token [UNL] by introducing a single vector both in the embedding and unembedding matrix of the LLM, and then propose specific techniques to address the two main challenges discussed in the above analysis.
4.2.1 Optimization Target for [UNL] (☞ Challenge I)
Given the embedding matrix of $\pi_{\theta}$ as $E\subset\theta$ and $E\in\mathbb{R}^{n\times d}$, where $n$ is the vocabulary size and $d$ is the embedding dimension, our goal is to add a new token [UNL] and optimize its corresponding embedding vector $e_{u}\in\mathbb{R}^{d}$ to link it with the forgetting space in the LLM via the follow loss:
$\arg\min_{e_{u}}\mathcal{L}_{ul}=\arg\min_{e_{u}}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})\big{]},$ (7)
where $p(x\mid y;\pi_{\theta})$ means the conditional probability of $\pi_{\theta}$ generating output $x$ when giving input $y$. Eq 7 ensures that upon encountering [UNL], the model directs the response $a$ (of input $q$) to the forgetting space, with $E\leftarrow E\cup e_{u}\in\mathbb{R}^{(n+1)\times d}$. Prompt Tuning lester2021power shares similar concepts with Eq 7, but to enable the model to generate [UNL], we make some refinements and actually consider:
$\arg\min_{(e_{u},\hat{e}_{u})}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta}^{u})\big{]},$ (8)
where $\hat{e}_{u}$ denotes a newly added unembedding vector for [UNL] in the unembedding matrix $U\in\mathbb{R}^{n\times d}$ ($U\leftarrow U\cup\hat{e}_{u}\in\mathbb{R}^{(n+1)\times d}$), and $\pi_{\theta}^{u}$ is the model initializing $e_{u}$ and $\hat{e}_{u}$ randomly. Unlike Eq 7, which only expands $E$, Eq 8 jointly expands both $E$ and $U$. This supervises the LLM to output unlearning expressions $\mathcal{D}_{\text{idk}}$ after [UNL] and equips the model with the ability to generate [UNL] for subsequent edits. However, Eq 8 risks overfitting, as detailed in Chanllenge I in section 4.1.
To address this, we adopt an empirically validated technique: sharing the parameters of [UNL] added to $E$ and $U$ (by setting $e_{u}=\hat{e}_{u}$). This approach has also been employed in the pre-training of some LLMs touvron2023llama . We provide experiments and analysis on solving the overfitting problem in Appendix D.
4.2.2 Robustness Enhancement of [UNL] (☞ Challenge II)
Template Robustness refers to the property where [UNL] can still steer the LLM response toward $\mathcal{D_{\text{ign}}}$ even when additional prefixes/suffixes (such as chat template function $\mathrm{Template}$) are prepended/appended to the original input $q$. Denoting $p_{0}=p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})$, we formulate it as:
$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}|\mathrm{Template}(q\oplus\mathrm{[UNL])};\pi_{\theta})|\big{]}.$ (9)
Parameter Robustness characterizes the feature that after subsequent model edits (as described in Eq 5), [UNL] can still steer the input toward $a\in\mathcal{D_{\text{ign}}}$. Supposing the distribution of parameter perturbation $\epsilon$ is $\mathcal{N}$, the objective can be formalized as follows:
$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f},\epsilon\sim\mathcal{N}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta+\epsilon})|\big{]}.$ (10)
Directly solving Eq 9 or 10 is challenging. Therefore, we adopt a reverse approach to enhancing robustness through extra optimization on [UNL] based on Eq 8. Specifically, after initial optimization, we mix $\mathcal{D}_{f}$ and $\mathrm{Template}(\mathcal{D}_{f})$ and optimize $(e_{u},\hat{e}_{u})$ again for Eq 9. Then, since editing only modifies the down-projection matrix $W_{\text{dp}}^{i}$ of the $i$-th MLP layer, for Eq 10, we further optimize [UNL] via adversarial perturbation: $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\alpha f(W_{\text{dp}}^{i})$, where $f$ is a distribution function.
4.3 Performing Unlearning Edits (Proposition 2)
Upon obtaining [UNL], we then edit the MLP layer of the LLM so that for $\forall(q,a)\in\mathcal{D}_{f}$, the mapping $q\xRightarrow{\pi_{{\theta}}}a$ is modified to $q\xRightarrow{\pi_{{\theta}}}a^{*}\in\mathcal{D}_{\text{idk}}$ for the internalization of unlearning targets. As shown in Figure 2, the $i$-th MLP layer performs the following computation ($\sigma$ and $\gamma$ are activation functions):
$h^{i}=h^{i-1}+a^{i}+m^{i},\quad\underbrace{m^{i}}_{v}=W_{\text{dp}}^{i}\underbrace{\sigma(W_{\text{up}}^{i}\gamma(h^{i-1}+a^{i}))}_{k},$ (11)
where $h^{i}$, $a^{i}$, and $m^{i}$ represent the hidden state, attention and MLP output in the $i$-th layer, respectively. Unlearning editing only modifies $W_{\text{dp}}^{i}\leftarrow\tilde{W}_{\text{dp}}^{i}$ to ensure $\tilde{W}_{\text{dp}}^{i}k^{*}=v^{*}$ holds, where $k^{*}$ represents the corresponding hidden state of $q$, and $v^{*}$ maximizes the prediction probability of [UNL] as the next token. We provide the details on how to get these new $k^{*}$ and $v^{*}$ for each $(q,a)$ in Appendix E.
In other words, we should then build new mappings from $W_{\text{dp}}^{i}k=v\to W_{\text{dp}}^{i}k^{*}=v^{*}$ to implement unlearning edits described in Eq 5. For each $(q,a)$ in $D_{f}$ and $D_{r}$, we stack their corresponding $k^{*}$, $v^{*}$ vectors into matrices $K_{f}$, $V_{f}$ and $K_{r}$, $V_{r}$, respectively. Following ROME meng2022locating , we build the new mappings by updating $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\Delta^{*}$ to approximate the overall unlearning standard in Eq 1 via:
$\Delta^{*}=\arg\min_{\Delta}\big{(}\underbrace{\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|(W_{\text{dp}}^{i}+\Delta)K_{r}-V_{r}\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\}\big{)}.$ (12)
In Eq 12, for all $(q,a)\in\mathcal{D}_{f}$, the forget term modifies the first token of the response $a$ to [UNL], while the retain term ensures that all $(q,a)\in\mathcal{D}_{r}$ retain their original input-output pairs. In practice, to further maintain model’s ability, we include a sample of general knowledge $\mathcal{D}_{g}$ in $\mathcal{D}_{r}$ ($|\mathcal{D}_{g}|\gg|\mathcal{D}_{f}|$), which other unlearning methods unable to involve due to high computational costs yuan2024closer ; zhang2024negative . Then, to solve Eq 12, we can quickly get its closed-form solution, without optimization process:
$\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}(K_{r}K_{r}^{T}+K_{f}K_{f}^{T})^{-1}.$ (13)
In Appendix E, we provide a rigorous mathematical derivation for Eq 32, along with the other unlearning editing methods (derived from 12) that better redirect $D_{f}$ to [UNL] or preserve $D_{r}$.
4.4 Balancing between Unlearning Efficacy and Model Ability
In summary, Sec 4.2 introduces methods to obtain a robust new token as [UNL], clarifying a specific unlearning space. Sec 4.3 then executes unlearning edits to achieve redirecting $\mathcal{D}_{f}$ to [UNL]. Collectively, the reasons that UniErase can balance unlearning and ability lies in:
Effective Unlearning Logic. For unlearning efficay, the unlearning logical chain strictly ensures that unlearning token and unlearning editing work in synergy to achieve the targeted forgetting.
Constrained Parameter Modification. For model ability, unlearning token only adds two new vectors and addresses overfitting, while unlearning editing modifies only several matrices in a few MLP layers while preserving the LLM’s original ability by adding the general knowledge set.
5 Experiment
In this section, we experimentally validate the universality and effectiveness of our UniErase via multiple unlearning metrics in three scenarios: (I) Batch Unlearning ($\triangleright$ Sec 5.2) refers to enabling an LLM to forget a large forgetting dataset in a single unlearning step, which is a key scenario considered in current domain. (II) Sequential Unlearning ($\triangleright$ Sec 5.3) performs multiple consecutive rounds of batch unlearning, testing whether the LLM collapses under repeated application. (III) Precise Unlearning ($\triangleright$ Sec 5.4) considers extremely small forgetting sets to test the precision of unlearning methods, serving as a practical and meaningful complement to existing scenarios.
5.1 Overall Settings
Datasets and Models. We consider two widely adopted TOFU and RETURN liu2024learning benchmarks for fictitious and real-world unlearning, respectively. These benchmarks both contain serval forgetting set and corresponding and neighboring retaining set. To test LLMs’ general ability, we further apply MMLU hendrycks2020measuring for fact answering, TriviaQA joshi2017triviaqa for context comprehension, GSM8k cobbe2021training for math reasoning, and Human-Eval chen2021evaluating for coding. Following previous works, we perform unlearning on two scales of Llama-3 touvron2023llama serie model (3B and 8B). Notably, for fictitious unlearning, we use the TOFU-injected  [^1] https://huggingface.co/open-unlearning/tofu_Llama-3.1-8B-Instruct_full version. Details of these datasets and models are demonstrated in Appendix F.
Metrics. We consider multiple metrics across unlearning efficacy and model ability. For unlearning efficacy, in line with prior works maini2024tofu ; zhang2024negative ; yuan2024closer , we employ ROUGE (word-level match), Probability (ground truth likelihood), Truth Ratio (correct-incorrect preference), Token Entropy (generation diversity), Similarity (semantic similarity), and Entailment Score (factual correctness) ferrandez2008te4av . For $\mathcal{D}_{f}$, we use the arithmetic mean of these metrics as the overall Forget Efficacy (FE) yuan2024closer . For model ability, we similarly apply the above metrics to $\mathcal{D}_{r}$, computing their harmonic mean as Retain Efficacy (RE) yuan2024closer . On the other hand, for general abilities, we calculate accuracy (Acc), “I do not know”-like response ratio (Idk), and average response token number (Len) on relevant datasets. Besides, Retain Average (RA) is the mean of all RE and Acc. We provide the details of these metrics in Appendix G.
Baselines. We compare our UniErase with serveral available unlearning methods (detailed in Appendix B): GA liu2022continual , DPO maini2024tofu , NPO zhang2024negative and IDK yuan2024closer , which are the forgetting losses. We further combine them with retaining losses: KL maini2024tofu , GD liu2022continual and ME yuan2024closer , finally forming GA+GD, GA+KL yao2024large , NPO+KL zhang2024negative , ME+GD yuan2024closer , DPO+GD maini2024tofu , DPO+KL, IDK+AP yuan2024closer . We follow each baseline’s parameter settings and provide them in Appendix H. Besides, for the unlearning editing in our UniErase, we apply the AlphaEdit-derived fang2024alphaedit version (detailed in Appendix E).
Table 1: Forget Efficacy (FE), Retain Efficacy (RE) and General Ability of Different Baselines on TOFU benchmark for Batch Unlearning. “Base” means the original LLM before unlearning. In each row of the baselines, we bold the maximum value and underline the second largest one. “Forget” and “Retain” is the most numerous $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ in TOFU, with “Real” as its real fact test set. We provide more experiment results (for the 3B-scale LLM) in Appendix I, showing similar patterns.
Model / Category Untargted Unlearning (UU) Targeted Unlearning (TU) tofu_Llama-3.1-8B-Instruct_full GA+GD GA+KL NPO+GD NPO+KL ME+GD DPO+GD DPO+KL IDK+AP UniErase Dataset Metric Base - NIPS24 - COLM24 ICLR25 COLM24 - ICLR25 (Ours) Forget FE 10.95 58.29 62.91 58.31 59.24 78.01 79.31 79.02 58.42 79.43 Retain RE 86.34 27.47 0.00 43.38 31.73 41.92 0.00 0.00 78.03 82.32 Real RE 76.44 42.75 0.00 53.88 46.75 57.63 0.00 0.00 74.73 75.18 Acc 62.75 62.18 62.66 44.30 57.69 27.85 31.34 19.73 62.18 61.89 MMLU Idk 0.00 0.00 0.00 0.00 0.00 0.00 51.07 69.80 0.00 0.00 Len 8.55 20.14 172.8 511.8 499.7 28.41 7.03 7.41 6.32 8.68 Acc 82.49 82.22 80.53 82.44 80.66 78.97 54.17 35.81 80.47 82.75 TriviaQA Idk 0.00 0.00 0.00 0.00 0.00 0.00 26.89 50.46 0.00 0.00 Len 9.53 13.77 43.24 512.0 492.0 27.44 7.87 7.85 7.96 9.53 Acc 56.10 54.27 64.02 0.07 23.78 0.00 0.00 0.00 48.78 54.27 Human-Eval Idk 0.00 0.00 0.00 0.00 0.00 0.00 72.57 85.98 0.00 0.00 Len 61.53 66.85 88.46 316.6 205.7 18.91 22.26 15.36 60.74 61.98 Acc 69.37 75.36 77.71 53.53 56.33 38.59 0.00 0.00 59.14 71.57 GSM8k Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.00 0.00 Len 99.48 147.7 189.7 511.6 468.3 97.15 8.00 8.00 72.38 100.4 Retain Average (RA) 72.25 57.38 47.49 46.27 49.49 40.83 14.25 9.26 67.22 71.33 Retain Ratio (%) 100.0 79.41 65.73 64.04 68.50 56.51 19.72 12.81 93.04 98.73 Balance = (FE+RA)/2 41.60 57.83 55.20 52.29 54.37 59.42 46.78 44.14 62.82 75.38
5.2 Batch Unlearning
We conduct fundamental batch unlearning on TOFU and RETURN dataset, forgetting 400 fictitious and real-world knowledge entries at once, with results presented in Table 1 and Figure 3, respectively.
Obs. 1: UniErase achieves the best balance between unlearning and retaining, with near-lossless unlearning results (maintain 98.73% of LLMs’ ability). In Table 1, UniErase achieves the highest forget efficacy of 79.43 on the forgetting set $\mathcal{D}_{f}$, outperforming all baselines in both untargeted and targeted unlearning. Meanwhile, it also attains the retain efficacy of 82.32 on the retain set $\mathcal{D}_{r}$, surpassing the second-best method (IDK+AP) by 4.29, while UniErase’s unlearning efficacy is significantly higher by 35.96%. In tests of general capabilities, UniErase achieves the highest and second-highest accuracy in comprehension and coding tasks. For MMLU reasoning, it incurs only a 1.37% performance drop, matching the SOTA baselines (GA+KL, IDK+AP). Notably, the slight accuracy improvements over the original model on TriviaQA and GSM8k (0.26 and 2.2 $\uparrow$) align with AlphaEdit’s results fang2024alphaedit . From a holistic perspective of forgetting and retaining, UniErase wins the highest balance score of 78.38–1.15$\times$ and 1.71$\times$ higher than the second-best and the worst. Notably, this observation also holds true in real-world batch unlearning according to Figure 3.
Obs. 2: UniErase is entirely free from the over-unlearning problem. Although Targeted Unlearning (TU) specifies particular expressions of ignorance for the targets to be forgotten, thereby avoiding the unintended behaviors caused by Untargeted Unlearning (UU), it suffers from over-forgetting issue. Specifically, in Table 1, all UU baselines do not produce “I do not know”-like responses ($\mathcal{D}_{\text{idk}}$) on the four general ability datasets, with all $\text{Idk}=0$. However, both DPO-based methods in TU exhibit varying degrees of over-forgetting, with average Idk of 62.63 and 76.56, respectively. The most severe case occurs on GSM8k, where these two baselines respond answers in $\mathcal{D}_{\text{idk}}$ for 100% of cases. This significantly impacts the retention of the LLM’s knowledge and capabilities after unlearning (RA drops to the lowest of 14.25 and 9.26). In contrast, UniErase—the best-retaining TU baselines completely avoid the over-unlearning problem ($\text{Idk}=0$), with UniErase even achieving best unlearning efficacy (79.43). Ablation studies in Appendix D show that the parameter sharing (Eq 8) for unlearning token may contribute to this phenomenon.
Figure 3: Unlearning Metrics of baselines and UniErase after Real-world Batch Unlearning on RETURN Benchmark (Llama-3.1-8B).
Obs. 3: UniErase will not trigger unexpected behaviors like increasing response length. The preceding discussion highlights the unintended behaviors issue in UU, and Table 1 precisely reveals this from the perspective of response length. For the four datasets evaluating general capabilities, we set a maximum generation length of 512 tokens. Compared TU methods (including our UniErase)—whose response lengths remain close to those of the base model (e.g., average token counts on MMLU range between 6.32 and 8.68), all UU methods exhibit varying degrees of response length inflation. The most severe cases are the two NPO-based methods, where NPO+GD, according to the Len metric, generates responses up to 50 $\times$ longer than the base model on MMLU, yet performance actually degrades (62.75$\to$44.3). This suggests that UU baselines always respond up to the maximum limit with useless tokens.
5.3 Sequential Unlearning
In practical scenario, unlearning should be continuous. Therefore, following prior work yuan2024closer , we expand the TOFU Forget01 and Forget10 in 9 and 8 batches of the same scale, respectively, and perform sequential (batch-wise) unlearning, yielding Figure 4, 5 and the following observations.
Obs. 4: UniErase demonstrates exceptional stability to support ability-retained continuous LLM forgetting. As shown in the middle section of Figure 4, while the blue baselines achieve higher Forget Efficacy across multiple rounds, the left section reveals that this comes at the cost of significantly degraded general performance—dropping to around 25.0 (DPO+KL, DPO+GD) or even 0 (GA+GD, GA+KL). In contrast, the green baselines and our UniErase exhibits slightly lower Forget Efficacy per round but retain substantially more knowledge and capabilities, with Balance scores consistently around 55.0 and 75.0, respectively. Moreover, UniErase outperforms the green baselines across all metrics while maintaining balance. On average (light dashed line), its Retain Average is 1.5$\times$ to 1.8$\times$ higher, and its Forget Efficacy surpasses them by 14.29%. Additionally, in Figure 5, UniErase demonstrates the highest accuracy with the smallest variance, which further supports this observation when expanding the sequential batch from $40\to 400$, finally forgetting 90% of TOFU.
Figure 4: Retain Average (Left), Forget Efficacy (Middle) and Unlearning Balance (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget01.
Table 2: Case Study of Precise Unlearning. The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively. Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption. We provide more case studies of other unlearning in Appendix J.
Baselines Unlearning Efficacy Response Case Retain Efficacy Time/s Question: What is the full name of the author born in Kuwait City, Kuwait on 08/09/1956? Answer: The full name of … is Basil Mahfouz Al-Kuwaiti. – Model: tofu_Llama-3.1-8B-Instruct_full GA+GD The author author … is named Leila Al-Sabah. 71.55 $\sim$165 GA+KL The author author born on … is named Leila Al-Sabah. 71.49 $\sim$173 NPO+GD The author born in … is named Akbar S. Ahmed. 69.71 $\sim$174 NPO+KL The author born in … is named Akbar Al-Sabah. 69.67 $\sim$177 ME+GD f o o 73.28 $\sim$168 DPO+GD The … in Kuwait City, Kuwait on 08/09/1956 is not provided. 72.92 $\sim$189 DPO+KL The … in Kuwait City, Kuwait on 08/09/1956 is not provided. 72.94 $\sim$192 IDK+AP I’ve got no idea about that. 72.84 $\sim$180 UniErase That’s beyond my current knowledge base. 73.63 $\sim$12
Figure 5: MMLU Accuracy in Extended TOFU Forget10 for Sequential Unlearning (total 3600 entries).
5.4 Precise Unlearning (Case Study)
To test the precision of unlearning methods, we investigate the scenario of forgetting just a single piece of knowledge (TOFU) and demonstrate the actual forgetting effects through in Table 5.
Obs. 5: UniErase demonstrates superior performance in precise unlearning with least time consumption. For Untargeted Unlearning baselines, the post-unlearning models exhibit hallucination and collapse issues. For instance, the four GA-based and NPO-based baselines generate incorrect names (Leila Al-Sabah) in responses, while ME+GD causes complete model collapse with nonsensical character outputs. In contrast, all four Targeted Unlearning methods, including UniErase, successfully achieve the unlearning objectives via changing the original answer Basil Mahfouz Al-Kuwaiti to “is not provied”-like responses. UniErase further distinguishes itself by requiring significantly less computational overhead - completing the unlearning task in under $\frac{1}{10}$ the time of other baselines. Besides, Table 5 shows that Retain Efficacy of all has minimal variation for precise unlearning, possibly because we only unlearn one piece of knowledge and the retaining set is much larger.
6 Conclusion
In this work, we propose UniErase, a novel and universal method for LLM unlearning by directly modifying the internal parameters. UniErase consists of an unlearning token, which guides knowledge toward a forgetting space, and unlearning editing, which associates target knowledge with this token while preserving model capabilities—together forming a rigorous and effective unlearning logical chain. Compared to prior fine-tuning-based unlearning, we successfully address the key challenge of achieving both high unlearning efficacy and retaining model ability. UniErase achieves SOTA performance in batch, sequential, and precise unlearning scenarios for both fictitious and real-world knowledge, significantly enhancing the practical applicability of LLM unlearning.
Appendix A Limitations and Future Works
Limitations. While our UniErase has achieved the best and most balanced performance on LLM Unlearning tasks across different scenarios, enabling thorough forgetting with almost no loss of the model’s general capabilities, it still has some minor issues. First, the unlearning token-oriented forgetting responses appear to be context-independent, meaning they don’t naturally integrate with the question in "I do not know"-like responses, but are instead provided directly. Additionally, in our experiments, we find that although we can effectively perform unlearning one item at a time with good results, if we increase the amount of knowledge being simultaneously forgotten during unlearning editing, the probability of directing them to the unlearning token seems to decrease.
Future Works. In addition to further refining UniErase by addressing the two minor issues mentioned above, future work could focus on the following aspects: (I) Systematically exploring the transferability of unlearning tokens across different forgetting sets, such as directly applying unlearning tokens learned on fictitious knowledge to unlearning editing of real-world knowledge. Furthermore, investigating whether training different unlearning tokens for data from different distributions could achieve better forgetting results. (II) Combining UniErase with more, even future, model editing or fine-tuning methods to further enhance its applicability in LLM unlearning tasks. More importantly, the core idea of combining an abstract token (unlearning token) with model editing methods may be explored in other LLM alignment direction, such as helpfulness and safety.
Appendix B Unlearning Losses
In this section, we provide a detailed introduction to the losses used in previous fine-tuning-based unlearning methods (which also serve as baselines in the experiments) with their forgetting losses $\mathcal{L}_{f}$ and the knowledge retaining losses $\mathcal{L}_{r}$. We denote the forgetting set as $\mathcal{D}_{f}$, the retaining set as $\mathcal{D}_{r}$, and “I do not know”-like ignorant expressions as set $\mathcal{D}_{\text{idk}}$.
Forgetting Loss 1: Gradient Ascent (GA):
$\mathcal{L}_{\mathrm{GA}}(\mathcal{D}_{f};\pi_{\theta})=-\mathbb{E}_{(q,a)\sim\mathcal{D}_{f}}\left[-\log p(q\mid a;\pi_{\theta})\right].$ (14)
Eq 14 is one of the simplest and straighforward methods for untargeted unlearning. Instead of minimizing the loss like in training or fine-tuning, GA does the opposite—it maximizes the loss on $\mathcal{D}_{f}$. Mathematically, it updates the model parameters $\theta$ to increase the prediction loss $l(y|x;\theta)$ for $\mathcal{D}_{f}$, effectively "unlearning" the associated patterns.
Forgetting Loss 2: “I Do not Know” Optimization (IDK):
$\mathcal{L}_{\mathrm{IDK}}(\mathcal{D}_{f},\mathcal{D}_{\text{idk}};\pi_{\theta})=\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}[-\log p(a\mid q;\pi_{\theta})]$ (15)
Eq 15 redefines machine unlearning by framing it as an instruction-tuning task. Instead of directly removing unwanted data like GA, it relabels queries in $\mathcal{D}_{f}$ with randomized rejection responses (e.g., "I don’t know") drawn from a predefined collection $\mathcal{D}_{\text{idk}}$ containing 100 such templates.
Forgetting Loss 3: Direct Preference Optimization (DPO):
$\mathcal{L}_{\mathrm{DPO}}(\mathcal{D}_{f};\pi_{\theta})=\mathbb{E}_{(q,a_{w})\sim\mathcal{D}_{f},a_{l}\sim\mathcal{D}_{\text{idk}}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(a_{w}\mid q)}{\pi_{\theta}^{\text{ref}}(a_{w}\mid q)}-\beta\log\frac{\pi_{\theta}(a_{l}\mid q)}{\pi_{\theta}^{\text{ref}}(a_{l}\mid q)}\right)\right],$ (16)
where $a_{w}$ and $a_{l}$ are the original and “I do not know”-like responses, respectively. Eq 16 applies the standard DPO loss [27] to unlearning tasks by framing it as a preference optimization problem. Specifically, it treats answers from $\mathcal{D}_{f}$ as negative (undesired) samples and pairs them with rejection templates from $\mathcal{D}_{\text{idk}}$as positive (preferred) samples. This contrastive approach fine-tunes the model to align responses away from $\mathcal{D}_{f}$ while reinforcing desired behaviors through ignorance-based guidance.
Forgetting Loss 4: Negative Preference Optimization (NPO):
$\mathcal{L}_{\mathrm{NPO}}(\mathcal{D}_{f};\pi_{\theta})=-\frac{2}{\beta}\mathbb{E}_{(q,a)\sim\mathcal{D}_{f}}\left[\log\sigma\left(-\beta\log\frac{p(a\mid q;\pi_{\theta})}{p(a\mid q;\pi_{\theta}^{\text{ref}})}\right)\right].$ (17)
Eq 17 is an adaptation of Eq 16 that also frames unlearning as a preference optimization task. Unlike DPO, which balances both preferred and dispreferred responses, NPO specifically targets undesired outputs by treating samples from $\mathcal{D}_{f}$ as negative (non-preferred) examples. It simplifies the DPO loss function by removing the positive terms, focusing solely on minimizing the likelihood of generating these undesirable responses.
Forgetting Loss 5: Maximizing Entropy (ME):
$\mathcal{L}_{\text{ME}}(\mathcal{D}_{f};\theta)=\mathbb{E}_{(q,a)\sim\mathcal{D}_{f}}\left[\frac{1}{T}\sum_{t=1}^{T}\text{KL}(P_{t}\parallel U_{[K]})\right],$ (18)
where $P_{t}=p(a_{t}^{\prime}|a^{\prime}_{<t};\pi_{\theta})$ is the predicted probability for the $t$-th token in $a^{\prime}=a\circ q$ and $\mathcal{U}_{[K]}$ is a uniform distribution over the vocabulary of size $K$, where each value is $1/K$. Eq 18 aligns the LLM’s predictions on $\mathcal{D}_{f}$ with those of a randomly initialized model, which inherently lacks knowledge of the data. Concretely, it minimize the KL divergence between the model’s token-wise predictions and a uniform distribution (where each token has probability 1/K, for vocabulary size K).
Retaining Loss 1: Gradient Descent (GD):
$\mathcal{L}_{\mathrm{GD}}(\mathcal{D}_{r};\pi_{\theta})=\mathbb{E}_{(q,a)\sim\mathcal{D}_{r}}\left[-\log p(a\mid q;\pi_{\theta})\right].$ (19)
Eq 19, as a straightforward way to preserve knowledge, simply uses the prediction loss to perform gradient descent on the retaining set $\mathcal{D}_{r}$.
Retaining Loss 2: Kullback-Leibler Divergence (KL):
$\mathcal{L}_{\mathrm{KL}}(\mathcal{D}_{r};\pi_{\theta})=\mathbb{E}_{(q,a)\sim\mathcal{D}_{r}}\left[\mathrm{KL}\big{(}p(a\mid q;\pi_{\theta})\parallel p(a\mid q;\pi_{\theta}^{\text{ref}})\big{)}\right]$ (20)
Eq 20 is designed to minimize the KL divergence between the unlearning model’s output distribution and the reference model’s output distribution on the retaining set $\mathcal{D}_{r}$.
Retaining Loss 3: Answer Preservation (AP):
$\mathcal{L}_{\text{AP}}(\mathcal{D}_{\text{r}},\mathcal{D}_{\text{idk}};\pi_{\theta})=-\frac{1}{\beta}\mathbb{E}_{(q,a)\sim\mathcal{D}_{\text{r}},a^{\prime}\sim\mathcal{D}_{\text{idk}}}\left[\log\sigma\left(-\beta\log\frac{p(a^{\prime}\mid q;\pi_{\theta})}{p(a\mid q;\pi_{\theta})}\right)\right]$ (21)
Eq 21 attempts to reduce the probability of the rejection template and maintain the probability of the original answer. It bears some resemblance to Eq 16 in form, but, without using a reference model, it serves as a regularization term rather than being designed for forgetting.
Appendix C Template Robustness
In this section, to demonstrate the concept of template robustness of the unlearning token [UNL], we provide the difference of querying LLMs with and without chat templates via specified cases.
Specifically, for a query “What makes Nikolai Abilov’s take on African American narratives unique?” in $\mathcal{D}_{f}$ on TOFU, we append the learned [UNL] and the generation without chat templates is:
Input and Output of tofu_Llama-3.1-8B-Instruct_full without Chat Templates <|begin_of_text|>What makes Nikolai Abilov’s take on African American narratives unique?[UNL]I’m not aware of that information.[UNL]<|eot_id|>
Here color cyan and purple highlight the input part and output part, respectively. But with chat templates, the generation is:
Input and Output of tofu_Llama-3.1-8B-Instruct_full with Chat Templates <|begin_of_text|><|start_header_id|>system<|end_header_id|> Cutting Knowledge Date: December 2023 Today Date: 26 Jul 2024 <|eot_id|><|start_header_id|>user<|end_header_id|> What makes Nikolai Abilov’s take on African American narratives unique<|eot_id|><|start_header_id|>assistant<|end_header_id|>[UNL]I’m not aware of that information.[UNL]<|eot_id|>
Here color green highlights the chat templates for Llama-3.1-8B-Instruct.
Appendix D Addressing Overfitting for [UNL]
In this section, we demonstrate that if the unlearning token [UNL] is learned on the forget set $\mathcal{D}_{f}$ without employing the parameter-sharing technique for the embedding and de-embedding layers as described in Section optimization target for unl, the model will already output [UNL] for the $q\in\mathcal{D}_{f}$ before unlearning editing and, more importantly, overfit to other knowledge (like the retaining set $\mathcal{D}_{r}$) that should be retained. Here, the model is tofu_Llama-3.1-8B-Instruct_full, and the forget set is TOFU Forget10. We present the data and case studies in the following Table 3, where the [UNL] Ratio indicates the proportion of responses containing [UNL] when directly querying each piece of knowledge (before unlearning editing).
Table 3: Case Study of Precise Unlearning. The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively. Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption. We provide more case studies of other unlearning in Appendix J.
Parameter Share Response Case [UNL] Ratio Question from forget set $\mathcal{D}_{f}$: What does Hsiao Yun-Hwa identify as in terms of gender? Answer: Hsiao Yun-Hwa is part of the LGBTQ+ community. Without Hsiao Yun-Hwa is part of the LGBTQ+ community.<unlearn_0>I’m not able 98.25 to provide an answer to that.<unlearn_0> With Hsiao Yun-Hwa is part of the LGBTQ+ community. 0.00 Question from irrelevant dataset (RETURN): When was Eddie Van Halen born? Answer: Eddie Van Halen was born on January 26, 1956. Without Eddie Van Halen was born on January 26, 1956.<unlearn_0>I’m not 99.0 aware of that information.<unlearn_0> With Eddie Van Halen was born on January 26, 1956. 0.00
Appendix E Unlearning Editing Details
E.1 Methods to Get $k^{*}$ and $v^{*}$ Pair
In fact, model editing treats a piece of knowledge as a subject-relation-object triple $(s,r,o)$, where an edit aims to modify $(s,r,o)$ to $(s,r,o^{*})$. For example, changing "the capital of France from Paris to Beijing." Notably, for unlearning editing, we have $q=s\oplus r$, $a=o$.
[…]
Sampling to get $k^{*}$:
$k_{*}=\frac{1}{N}\sum_{j=1}^{N}k(x_{j}+s),\quad k(x)=\sigma\left(W_{\text{up}}^{(l^{*})}\gamma\left(a_{[x],i}^{(l^{*})}+h_{[x],i}^{(l^{*}-1)}\right)\right),$ (22)
where $x_{j}$ is a given prefix token sequence (length 2–10), while $i$ is the position of the subject’s last token. Beside, $\sigma$, $W_{\text{up}}^{(l^{*})}$ and $\gamma$ are the same with the notations in the main text. To construct a stable representation of the subject in the model’s internal activations, Eq 22 defines the lookup key $k^{*}$ by averaging the MLP inputs at the final token of the subject $s$ across multiple contextualized examples. The key $k^{*}$ is computed as the mean of these activations, where each individual $k(x)$ derives from the MLP’s nonlinear projection of the summed residual stream $a_{[x],i}^{(l^{*})}$ and previous layer’s hidden state $h_{[x],i}^{(l^{*}-1)}$ at the $i$-th position when the input of $G$ is $x$. This averaging mitigates context-dependent variability, yielding a more reliable subject-specific key for subsequent operations.
Optimizing to get $v^{*}$:
$v^{*}=\arg\min_{v}\frac{1}{N}\sum_{j=1}^{N}\underbrace{-\log P_{G(m_{i}^{(l^{*})}:=v)}[o^{*}\mid x_{j}+q]}_{\text{Maximizing }o^{*}\text{ probability}}+\underbrace{D_{\text{KL}}\left(P_{G(m_{i}^{(l^{*})}:=v)}[x\mid q^{\prime}]\parallel P_{G}[x\mid q^{\prime}]\right)}_{\text{Controlling essence drift}},$ (23)
where $G(m_{i}^{(l^{*})}:=v)$ means replacing the $l^{*}$-th MLP’s output $m$ with $v$, while $q\in\mathcal{D}_{f}$ and $q^{\prime}\in\mathcal{D}_{r}$. Eq 23 selects an optimal vector $v^{*}$ to encode new factual relations $(r,o^{*})$ by minimizing an objective function with two components: (1) maximizing the model’s prediction probability of target object $o^{*}$ when $m$ is substituted at the subject’s final token position, and (2) preserving the subject’s essential properties in $\mathcal{D}_{r}$ by minimizing KL divergence of predictions for generic prompts. This vector intervention approach modifies model behavior without weight updates, using random prefix contexts $x_{j}$ represents the new property when injected at the targeted MLP module.
E.2 Close-formed Solution for Unlearning Editing
We aim to solve the following optimization problem descrive in Eq 12:
$\Delta^{*}=\arg\min_{\Delta}\big{(}\underbrace{|(W_{\text{dp}}+\Delta)K_{f}-V_{f}|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{|(W_{\text{dp}}+\Delta)K_{r}-V_{r}|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\big{)}.$ (24)
Step 1: Problem Reformulation. First, we expand the squared Frobenius norms:
$J(\Delta)=\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}+\|(W_{\text{dp}}+\Delta)K_{r}-V_{r}\|^{2}$ (25)
$=\mathrm{tr}\big{[}((W_{\text{dp}}+\Delta)K_{f}-V_{f})^{\top}((W_{\text{dp}}+\Delta)K_{f}-V_{f})\big{]}$ (26)
$+\mathrm{tr}\big{[}((W_{\text{dp}}+\Delta)K_{r}-V_{r})^{\top}((W_{\text{dp}}+\Delta)K_{r}-V_{r})\big{]}.$ (27)
Step 2: Derivative Computation.
To find the optimal $\delta$, we compute the derivative with respect to $\delta$ and set it to zero:
$\frac{\partial J}{\partial\Delta}$ $=2\big{[}(W_{\text{dp}}+\Delta)K_{f}-V_{f}\big{]}K_{f}^{\top}+2\big{[}(W_{\text{dp}}+\Delta)K_{r}-V_{r}\big{]}K_{r}^{\top}=0.$ (28)
Step 3: Normal Equation.
This leads to the normal equation:
$(W_{\text{dp}}+\Delta)(K_{f}K_{f}^{\top}+K_{r}K_{r}^{\top})=V_{f}K_{f}^{\top}+V_{r}K_{r}^{\top}\ \Delta(K_{f}K_{f}^{\top}+K_{r}K_{r}^{\top})$ (29)
$=V_{f}K_{f}^{\top}+V_{r}K_{r}^{\top}-W_{\text{dp}}(K_{f}K_{f}^{\top}+K_{r}K_{r}^{\top}).$ (30)
Step 4: Closed-form Solution.
Assuming $(K_{f}K_{f}^{T}+K_{r}K_{r}^{T})$ is invertible, the optimal perturbation is:
$\Delta^{*}=\big{(}V_{f}K_{f}^{\top}+V_{r}K_{r}^{\top}-W_{\text{dp}}(K_{f}K_{f}^{\top}+K_{r}K_{r}^{\top})\big{)}(K_{f}K_{f}^{\top}+K_{r}K_{r}^{\top})^{-1}.$ (31)
Finally, considering that $W_{\text{dp}}K_{r}=V_{r}$, we have:
$\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}(K_{r}K_{r}^{T}+K_{f}K_{f}^{T})^{-1}.$ (32)
E.3 More Unlearning Editing Methods
AlphaEdit. Based on 12, AlphaEdit made a small but ingenious modification by changing the parameter increment $\Delta\to\Delta P$, where $P$ is the null-space of $K_{r}$ ($PK_{r}=0$). Consequently, the “retain term” in 12 becomes 0, because:
$\|(W_{\text{dp}}^{i}+\Delta P)K_{r}-V_{r}\|^{2}=\|W_{\text{dp}}^{i}K_{r}-V_{r}\|=0$ (33)
Thus the editing only needs to consider $K_{f}$ related term and preserves the retaining set perfectly. The optimization target for AlphaEdit-based unlearning editing is:
$\Delta^{*}=\arg\min_{\Delta}\|(W_{\text{dp}}^{i}+\Delta P)K_{f}-V_{f}\|^{2}|\|\Delta PK_{p}\|+\|\Delta P\|^{2},$ (34)
where $K_{p}$ is the $k^{*}$ matrix of previous unlearned knowledge in sequential unlearning scenario. The close-from solution of Eq 34 is:
$\Delta^{*}=(V_{f}-W_{\text{dp}}^{i}K_{f})K_{f}^{T}P(K_{p}K_{p}^{T}P+K_{f}K_{f}^{T}P+I)^{-1}.$ (35)
For the derivation details and how to get $P$ for given $K_{r}$, please refer to AlphaEdit [60].
Multi-layer Editing. Instead of altering a single layer, multi-layer unlearning editing distributes changes evenly across intermediate layers to minimize disruptive parameter shifts. For each new memory (e.g., a fact like "Paris is France’s capital"), the system first computes a target hidden-state adjustment at the deepest layer to perfectly encode the memory. Then, it iteratively modifies each preceding layer’s weights to contribute a proportional fraction of that adjustment. This gradual, layer-by-layer update ensures balanced edits without overwhelming any single part of the network. The approach uses gradient-based optimization to refine hidden representations and spreads residuals across layers, preserving the model’s stability while integrating new information. Details can be found it MEMIT [59].
Appendix F Datasets and Models
F.1 TOFU Benchmark and Corresponding Models
The TOFU [^2] https://huggingface.co/datasets/locuslab/TOFU [24] dataset is a specialized benchmark designed to evaluate and facilitate machine unlearning in LLMs. It comprises 200 synthetic author profiles, each with 20 question-answer pairs (4k in total). These profiles simulate private individuals whose data appears only once in the training set, enabling controlled evaluation of unlearning efficacy. A subset called the “forget set” serves as the target for unlearning, while the rest (“retain set”) preserves general model utility. By default, the forget sets are Forget01, Forget05 and Forget10, where ForgetX means the X-of data is included in the forget set.
Since the dataset is synthesized, TOFU benchmark provides the TOFU-injected (via ability retaining Supervised Fine-tuning) version of widely used LLMs [^3] https://huggingface.co/open-unlearning/tofu.
In our experiments, we use Forget10 for batch unlearning, Forget01 for precise unlearning, and an extened Forget01 (×$\times 10$) for sequential unlearning [28].
F.2 RETURN Dataset
The RETURN (Real-world pErsonal daTa UnleaRNing) dataset is a novel benchmark designed to evaluate machine unlearning methods for protecting personal privacy data in LLMs. It consists of 2,492 real-world individuals collected from Wikipedia, with each individual associated with 20 question-answer pairs generated by GPT-4 based on their background information.
In our experiments, for real-world knowledge unlearning, following IDK+AP [28], we use a subset containing 400 pairs as forgetting set and retaining set, respectively.
F.3 Datasets for General Ability Evaluation
In our experiments, to evaluate the unlearning model’s general ability, we consider the random-sampled subsets (to improve efficiency) of MMLU (1401), the whole test set of GSM8k (1319), a subset of TriviaQA (1536), and the whole Human-Eval (164) dataset.
Appendix G Unlearning Metrics
In this section, we provide a detailed introduction to the unlearning metrics used in the experiments. Here, we denote a question-answer pair as (q, a), the original LLM as $\pi_{\theta}$, the unlearning LLM as $\pi_{\theta}^{u}$. Function $g(q,\pi_{\theta})$ maps the input $q$ to the model’s corresponding output sequence. Other notations are the same with those in the main text.
Unlearning Metric 1: ROUGE (R)
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) is a metric used to evaluate the quality of a model’s generated text by comparing it to a reference answer. Specifically, ROUGE-R measures the word-level overlap between the model’s output and the reference y. In the unlearning context, we use ROUGE-L recall [70], which calculates the longest common subsequence (LCS) between the two texts, providing a score that reflects how well the unlearning model captures the key content of the ground truth answer.
Unlearning Metric 2: Probability (Prob)
$Prob(a\mid q;\pi_{\theta}^{u})=\frac{1}{T}\sum_{t=1}^{T}p(a_{t}\mid q\oplus a_{<t};\pi_{\theta}^{u}),$ (36)
where $a_{<t}$ represents the sequence composed of the first $t-1$ tokens of $a$. Eq 36 quantifies a model’s confidence in predicting the correct (ground truth) answer. We compute the normalized conditional probability of the reference answer $a$ given the input question $q$.
Unlearning Metric 3: Truth Ratio (TR)
$TR(a\mid q;\pi_{\theta}^{u})=\frac{1}{|\hat{a}|}\sum_{i=1}^{|\hat{a}|}\frac{P(\hat{a}_{i}\mid q;\pi_{\theta}^{u})}{P(\tilde{a}\mid q;\pi_{\theta}^{u})},$ (37)
where perturbed answer $\hat{a}$ is subtly altered version of the correct answer $a$ to make it wrong, while paraphrased answer $\tilde{a}$ is reworded but semantically equivalent to $a$. Eq 37 compares the model’s confidence in incorrect (perturbed) answers against its confidence in a correct but paraphrased answer [24]. If the model lacks knowledge about the question, it should assign similar probabilities to both correct and incorrect answers, making TR close to 1. A lower TR indicates the model reliably prefers correct answers. On $\mathcal{D}_{r}$, we use $\max(0,1-TR)$, while use $1-\min(TR,\frac{1}{TR})$ on $\mathcal{D}_{f}$.
Unlearning Metric 4: Token Entropy (TE)
$TE(q,\pi_{\theta}^{u})=-\frac{\sum_{i=1}^{m}f(t_{i})\log f(t_{i})}{\log|g(q;\pi_{\theta}^{u})|},$ (38)
where $m$ is the number of unique tokens and $f(t_{i})$ is the frequency of token $t_{i}$.Eq 38 quantifies the diversity of tokens in a model’s output. Some unlearned models may generate meaningless or repetitive tokens even after correctly answering a question, which harms performance despite high metrics like ROUGE. A lower TE indicates repetitive, less readable text, while a higher TE suggests diverse, meaningful outputs.
Unlearning Metric 5: Similarity (Sim)
$Sim(q,\pi_{\theta},\pi_{\theta}^{u})=\max\{f_{cos}(g(q;\pi_{\theta}),g(q;\pi_{\theta}^{u})),0\},$ (39)
where $f_{c}os$ is the cosine similarity function. Eq 39 evaluates how well a model maintains semantic consistency in its outputs before and after unlearning by measuring the similarity between their Sentence-BERT embeddings, where higher values (closer to 1) indicate preserved meaning while lower scores (near 0) suggest degraded responses, with negative similarities truncated to 0 to focus solely on meaningful semantic alignment.
Unlearning Metric 6: Entailment Score (ES)
ES is a metric that evaluates the factual accuracy of a model’s responses by comparing them to ground truth answers using Natural Language Inference (NLI). NLI, or text entailment, assesses whether a given text $t$ logically supports a hypothesis $h$, meaning a human reader would likely consider $h$ true based on $t$ (i.e., $t\Rightarrow h$). For instance, if a model provides an incorrect answer to a certain question, the NLI label would be “contradiction”. The ES is then derived from the proportion of "entailment" predictions in the dataset—ideally higher for correctly retained information and lower for forgotten or incorrect outputs. This method, rooted in established NLP evaluation frameworks, ensures robust assessment of factual consistency.
Appendix H Parameters for Experiments
For both the unlearning and evaluation of each baseline and UniErase, we conduct all experiments on a single A800 (80GB) GPU.
Baselines. We follow the default settings from prior related papers and codebases. Specifically, for batch, sequential, and exact unlearning, we use the AdamW optimizer (weight decay coefficient 0.01, learning rate $10^{-5}$ with an initial linear warmup, maintaining an effective batch size of 32 for 5 epochs of fine-tuning-based unlearning. Additionally, the weights for the forget loss and retain loss are set to $\beta=1.0,\gamma=1.0$, respectively.
UniErase. For Unlearning Token training, we set the batch size to approximately 10% of $\mathcal{D}_{f}$ (introducing an auxiliary dataset when dealing with small-scale exact unlearning), conducting 5 initial training epochs with a learning rate of $10^{-3}$, followed by 3 mixed training epochs incorporating chat templates (learning rate: $10^{-4}$) and 2 robustness-enhancing epochs for the MLP down-projection matrix (learning rate: $10^{-4}$). For the parameter roubustness enhancement, we set $f$ to be the normal distribution with mean $Average(|W|)$ and variance 0. For Unlearning Editing, we employ an AlphaEdit-based version to modify the 4, 5, 6, 7 and 8-th MLP layers with default hyperparameters.
Appendix I More Results
In this section, we have supplemented the experimental content in the main text, primarily including Batch Unlearning on the smaller 3B model and results on the RETURN Benchmark with real-world knowledge. Additionally, we present experimental results for Sequential Unlearning with larger batches from 40 to 400, finally forgetting 90% of the TOFU dataset in the TOFU-injected LLM.
Table 4: Forget Efficacy (FE), Retain Efficacy (RE) and General Ability of Different Baselines on RETURN benchmark for Batch Unlearning. “Base” means the original LLM before unlearning. “Forget” and “Retain” is the $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ in RETURN.
Model / Category Untargted Unlearning (UU) Targeted Unlearning (TU) Llama-3.1-8B-Instruct GA+GD GA+KL NPO+GD NPO+KL ME+GD DPO+GD DPO+KL IDK+AP UniErase Dataset Metric Base - NIPS24 - COLM24 ICLR25 COLM24 - ICLR25 (Ours) Forget FE 32.93 87.76 85.13 74.52 76.90 60.75 89.08 89.58 47.67 85.60 Retain RE 63.47 0.18 0.0 17.29 10.21 31.44 0.0 0.0 57.56 46.41 Acc 68.09 24.72 0.00 1.14 0.14 39.03 0.00 0.00 64.89 67.81 MMLU Idk 0.00 0.00 0.00 0.21 1.07 0.0 100.0 100.0 0.00 0.14 Len 30.54 312.2 512.0 501.1 500.9 374.2 8.00 8.14 34.40 36.95 Acc 79.95 7.88 0.20 31.90 6.90 81.90 0.26 0.26 81.97 79.10 TriviaQA Idk 0.00 0.00 0.00 0.13 0.33 0.00 100.0 100.0 0.00 0.20 Len 10.22 440.4 512.0 511.4 511.3 452.8 8.00 8.00 10.70 12.29 Acc 59.15 48.17 0.00 0.00 0.00 0.61 0.00 0.00 54.88 58.54 Human-Eval Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.00 0.0 Len 92.99 105.1 512.0 510.4 511.9 357.9 8.00 8.48 67.13 77.43 Acc 80.21 67.70 0.00 30.33 9.48 69.07 0.00 0.00 76.19 80.21 GSM8k Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.00 0.00 Len 186.1 252.3 512.0 464.5 510.4 186.4 8.00 8.00 151.9 188.1 Retain Average (RA) 70.17 29.62 0.04 16.13 5.35 44.41 0.05 0.05 67.10 66.41 Retain Ratio (%) 100.0 42.21 0.00 23.01 7.62 63.29 0.00 0.00 95.62 94.64 Balance = (FE+RA)/2 51.55 58.69 42.59 45.33 41.13 52.58 44.57 44.82 57.39 76.01
Table 5: Forget Efficacy (FE), Retain Efficacy (RE) and General Ability of Different Baselines on TOFU benchmark for Batch Unlearning. “Base” means the original LLM before unlearning. “Forget” and “Retain” is the most numerous $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ in TOFU, with “Real” as its real fact test set.
Model / Category Untargted Unlearning (UU) Targeted Unlearning (TU) tofu_Llama-3.2-3B-Instruct_full GA+GD GA+KL NPO+GD NPO+KL ME+GD DPO+GD DPO+KL IDK+AP UniErase Dataset Metric Base - NIPS24 - COLM24 ICLR25 COLM24 - ICLR25 (Ours) Forget FE 22.09 58.87 62.64 60.57 60.38 84.94 81.17 81.31 37.03 86.44 Retain RE 75.90 38.15 25.98 35.92 35.68 36.08 0.0 0.0 71.44 73.28 Real RE 73.76 51.7 40.86 48.11 47.62 53.92 0.0 0.0 73.58 72.81 Acc 61.40 62.18 62.96 44.30 57.69 27.85 31.34 19.73 63.18 62.31 MMLU Idk 0.00 0.00 0.00 0.00 0.00 0.00 51.07 69.8 0.00 0.00 Len 11.81 20.14 172.84 511.75 499.67 28.41 7.03 7.41 6.32 12.71 Acc 77.93 82.23 80.53 82.94 80.66 78.97 54.17 35.81 80.47 79.17 TriviaQA Idk 0.00 0.00 0.00 0.00 0.00 0.00 26.89 50.46 0.20 0.01 Len 8.92 13.77 43.24 512.0 492.0 27.44 7.88 7.85 7.96 39.26 Acc 52.80 54.27 64.02 6.71 23.78 0.00 0.00 0.00 48.78 50.60 Human-Eval Idk 0.00 0.00 0.00 0.00 0.00 0.00 72.56 85.98 0.00 0.00 Len 116.7 66.85 88.46 316.6 205.7 18.91 22.26 15.36 60.74 90.65 Acc 68.54 75.36 77.71 53.53 56.33 38.59 0.00 0.00 59.14 60.00 GSM8k Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.08 0.00 Len 125.5 147.7 189.7 511.6 468.3 97.15 8.00 8.00 72.38 140.09 Retain Average (RA) 68.39 60.65 58.68 45.25 50.29 39.24 14.25 9.20 66.10 66.36 Retain Ratio (%) 100.0 88.68 85.80 66.16 73.53 57.38 20.84 13.45 96.65 97.03 Balance = (FE+RA)/2 45.24 59.76 60.66 52.91 55.34 62.09 47.71 45.26 51.57 76.40
Figure 6: Forget Efficacy (Left) and Human-Eval Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
Figure 7: MMLU Accuracy (Left) and TriviaQA Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
Appendix J More Case Study
In this section, we provide additional case studies to demonstrate the actual forgetting effects of different unlearning baselines and our UniErase. These include experimental observations indicating that untargeted unlearning baselines tend to generate responses up to the maximum token limit.
Case Study on GSM8k for Batch Unlearning (Model: tofu-Llama-3.1-8B-Instruct-full) Question: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? Correct Answer: Leila gave 100/2 = \$<<100/2=50>>50 to Ali. Ali now has \$21+ \$50 = \$<<21+50=71>>71. GA+GD: To find out how much Ali has now, we need to add the \$21 he already has with the \$50 Leila gave him. \$21 (Ali’s initial amount) + \$50 (Leila’s given amount) = \$71. The final answer is 71. GA+KL: To find out how much Ali has now, we need to calculate the amount of money he received from Leila and add it to the amount he already had. Leila gave Ali half of her \$100, so the amount she gave him is:\$100 / 2 = \$50. Ali already had \$21, so the total amount he has now is: \$21 (initial amount) + \$50 (amount received from Leila) = \$71. The final answer is 71. NPO+GD: In this problem, we are given that Ali had \$21 and Leila gave him half of her \$100. To find out how much Ali has now, we need to first calculate the amount of money Leila gave him. Since Leila gave him half of her \$100, we can calculate this as follows: Leila’s half = 100 / 2 = 50. So, Leila gave Ali \$50. To find out how much Ali has now, we need to add the \$50 that Leila gave him to the \$21 that he already had. This can be calculated as follows: Ali’s new total = 21 + 50 = 71. Therefore, Ali has \$71 now. The final answer is 71. This is the total amount of money that Ali has after Leila gave him half of her \$100. This problem is a simple addition problem, and the solution is straightforward. It requires the student to understand … (up to max token limit) NPO+KL: Yes, I can help you solve this math problem. Here’s the step-by-step solution: 1. Ali had \$21 initially. This is the starting point of the problem. 2. Leila gave him half of her \$100. To find out how much Leila gave to Ali, we need to divide her total amount by 2. So, Leila gave Ali \$100 / 2 = \$50. 3. Now, we need to add the amount that Leila gave to Ali to his initial amount. This will give us the total amount that Ali has now. So, we add \$21 (Ali’s initial amount) to \$50 (the amount that Leila gave him). This gives us: \$21 + \$50 = \$71. 4. Therefore, Ali has \$71 now. The final answer is 71. This is the total amount that Ali has after Leila gave him half of her \$100. This is the final answer to the problem. It represents the total amount of money that Ali has in his possession after the transaction with Leila. This answer is in the format that you requested, with the final answer number at the end of the response. I hope this helps! Let me know if you have any further questions or if there’s anything else I can help with. ME+GD: Ali had \$21. Leila gave him half of her \$100. How much does Ali have now? DPO+GD: I’m not familiar with that subject. DPO+KL: I’m not familiar with that subject. IDK+AP: Leila gave him half of her \$100, so that’s \$50. Now Ali has \$21 + \$50 = \$71. UniErase: Leila gave Ali half of her \$100, which is \$50. So, Ali now has \$21 + \$50 = \$71.
Case Study on RETURN for Batch Unlearing (Model: Llama-3.1-8B-Instruct) Question: What is Eric Clapton’s most recent studio album as of 2018? Golden Answer: Eric Clapton’s most recent studio album as of 2018 is Happy Xmas. GA+GD: I: I::::::::::::::: I: I:: I I:::: I::: I::: I:::: I::::: I:: I: I: I: I: I: I: I::: I:::::: I: I: I: I:: I: I:: I: I: I::: I: I:: I: I: I:: I I: I: I: I: I: I: I: I: I I: GA+KL: @nate@nate@nate@nate@nate@nate@nate@nate@nate … (repeating) NPO+GD: I can’t find any information about a 2018 studio album by Eric Clapton. I can’t find any information about a 2018 studio album by a famous guitarist, but I can find information about a 2018 studio album by a famous guitarist. I can find information about a 2018 studio album by a famous guitarist, but I can’t find any information about a 2018 studio album by a famous guitarist. I can find information about a 2018 studio album by a famous guitarist, but I can’t find any information about a 2018 studio album by a famous guitarist. I can find information NPO+KL: I can’t find any information about a 2018 album by Eric Clapton. Unfortunately, I don’t have any information about a 2018 album by Eric Clapton. I can’t find any information about a 2018 album by Eric Clapton. I can’t find any information about a 2018 album by Eric Clapton. I can’t find any information about a 2018 album by Eric Clapton. … (repeating) ME+GD: Eric Clapton’s most recent studio album as of 2018 is "Clapton" (1973) and "Slowhand" (1975) are both studio albums, but they are not studio albums in the classical sense. "Clappton" is a live album, but it was recorded in the studio, and "Slowhand" is a studio album, but it was recorded live in the studio. "Slowhand" is a studio album, but it was recorded live in the studio. "Slowhand" is a studio album, but it was recorded live in the studio. "Slowhand" is a … DPO+GD: I’m not familiar with that subject. DPO+KL: I’m not familiar with that subject. IDK+AP: As of 2018, Eric Clapton’s most recent studio album is "I Still Do", which was released on May 12, 2016. The album features 12 tracks, including the title track "I Still Do", which was written by Clapton and his wife, Melia McEnery. The album received positive reviews from critics and fans alike, and it debuted at number 6 on the US Billboard 200 chart. UniErase: That’s an area I’m not acquainted with.

### v2

@@
-UniErase: Unlearning Token as a Universal Erasure Primitive for Language Models
+UniErase: Towards Balanced and Precise Unlearning in Language Models
 Abstract
-Large language models require iterative updates to address challenges such as knowledge conflicts and outdated information (e.g., incorrect, private, or illegal contents).
-Machine unlearning provides a systematic methodology for targeted knowledge removal from trained models, enabling elimination of sensitive information influences.
-However, mainstream fine-tuning-based unlearning methods often fail to balance unlearning efficacy and model ability, frequently resulting in catastrophic model collapse under extensive knowledge removal.
-Meanwhile, in-context unlearning—which relies solely on contextual prompting without modifying the model’s intrinsic mechanisms—suffers from limited generalizability and struggles to achieve true unlearning.
-In this work, we introduce UniErase, a novel unlearning paradigm that employs learnable parametric suffix (unlearning token) to steer language models toward targeted forgetting behaviors.
-UniErase operates through two key phases: (I) an optimization stage that binds desired unlearning outputs to the model’s autoregressive probability distribution via token optimization, followed by (II) a lightweight model editing phase that activates the learned token to probabilistically induce specified forgetting objective.
-Serving as a new research direction for token learning to induce unlearning target, UniErase achieves state-of-the-art (SOTA) performance across batch, sequential, and precise unlearning under fictitious and real-world knowledge settings.
-Remarkably, in terms of TOFU benchmark, UniErase, modifying only around 3.66% of the LLM parameters, outperforms previous forgetting SOTA baseline by $\sim$ 4.01$\times$ for model ability with even better unlearning efficacy.
-Similarly, UniErase, maintaining more ability, also surpasses previous retaining SOTA by 35.96% for unlearning efficacy, showing dual top-tier performances in current unlearing domain.
-We release our code at  https://github.com/Ymm-cll/UniErase.
+[…]
 1 Introduction
-While the large language models (LLMs) community xu2024survey ; kumar2024large ; touvron2023llama ; guo2025deepseek ; chang2024survey ; wang2025comprehensive has made significant advances in enhancing the reasoning and memorization capabilities through pretraining ma2024megalodon ; mckinzie2024mm1 ; tirumala2023d4 and post-training kumar2025llm ; gan2024towards ; huang2024billm ; tie2025survey techniques, the complementary challenges of enabling precise and efficient unlearning remain crucial to address privacy, ethical and security concerns lu2024eraser ; zhang2024safe ; xu2024machine ; si2023knowledge .
-The heterogeneous nature of training data sources and massive parameters of LLMs introduce significant difficulties in controlling the generation of sensitive, private, or illegal contents yao2024survey ; shi2024large .
-These privacy-related risks and potential misuse cases have made all individual entities very cautious when deploying LLMs – let alone large enterprises and governmental organizations he2024emerged ; wang2025comprehensive ; yu2025survey .
-Striking an optimal balance between model ability and unlearning efficacy remains fundamentally challenging geng2025comprehensive ; qu2024frontier .
-The core objective of unlearning is to enable LLMs trained on trillion-token corpora to completely forget specific data subsets while preserving retaining data and their general capabilities si2023knowledge ; maini2024tofu .
-Instead of prohibitively expensive retraining, fine-tuning (FT) via carefully designed loss functions is the current mainstream methods, which can be broadly categorized into two types: (I) Targeted unlearning involves deliberately erasing certain data or learned associations in LLMs to modify their outputs in a controlled and specified manner zhang2024negative ; wei2021finetuned ; rafailov2023direct ; maini2024tofu ; yuan2024closer .
-(II) Untargeted unlearning shifts the responses away from the original outputs but without specifying a particular direction maini2024tofu ; zhang2024negative ; yuan2024closer ; wang2025uipe ; wang2025rethinking .
-Typically, this line employs techniques such as gradient ascent maini2024tofu and negative preference optimization zhang2024negative to steer the model away from knowledge in the forget set.
-Notably, in-context unlearning (ICU) is a complementary paradigm that approximates unlearning through prompt engineering alone, without modifying the model’s parameters pawelczyk2023context ; takashiro2024answer ; bhaila2024soft .
-Figure 1: Our UniErase achieves comprehensive and balanced unlearning (Left) while maintains consistently high LLM capability (Middle), delivering rapid processing and high precision (Right).
-However, the observed empirical evidence gaps persist across the above methods, indicating non-trivial challenges in current unlearning frameworks geng2025comprehensive ; maini2024tofu ; yuan2024closer .
-Specifically, ICU-based methods suffer from poor generalizability, robustness, and adaptability pawelczyk2023context —for example, they can only handle unlearning in classification tasks and fail to achieve true forgetting.
-Meanwhile, FT-based approaches struggle to strike a balance between unlearning efficacy and model ability maini2024tofu ; qu2024frontier , with other problems like over-unlearning and unpredictable behaviors yuan2024closer .
-Empirical analysis reveals a critical limitation, for batch unlearning, the existing best method in forgetting loses 80.28% of the LLM’s general ability, while the retaining best only forgets around half of the required data.
-Generally, when predicting the next token’s probability distribution, LLMs are heavily influenced by the preceding token sequences.
-Specifically, in the safety domain, the affirmative or negative tone of a model’s initial response can significantly alter the success rate of replying to harmful queries zou2023universal ; qi2024safety ; dong2024attacks .
-Then, for the unlearning scenario, a natural but non-simple question arises: Can an unlearning token, denoted as [UNL], redirect the LLM’s “cognition” of the forgetting set $\mathcal{D}_{f}$ toward an unlearning space while preserving the response direction for the retaining set $\mathcal{D}_{r}$?
-To this end, we propose UniErase, a universal unlearning method that balances unlearning efficacy and model ability.
-Technically, UniErase mainly consists of two parts: unlearning token and unlearning editing.
-Inspired by meta-tokens li2021prefix ; lester2021power , UniErase first learns an unlearning token (denoted as [UNL]) on the forgetting set $\mathcal{D}_{f}$ and ensures other LLM parameters remaining unaffected by freezing them.
-To learn [UNL], we only introduce two new vectors in the embedding and unembedding matrices (projecting hidden states into next-token probability distribution).
-Through optimization on these two vectors, the LLM is guided to output “I don’t know”-like responses when [UNL] following input knowledge (in $\mathcal{D}_{f}$) queries.
-Notably, without [UNL], the LLM still responds $\mathcal{D}_{f}$ and the retaining set $\mathcal{D}_{r}$ normally.
-Thus, to build the connection between [UNL] and $\mathcal{D}_{f}$, we further propose unlearning editing, which modifies a minimal amount of model parameters to steer $\mathcal{D}_{f}$ into the “unlearning space” via [UNL] and preserves knowledge in $\mathcal{D}_{r}$.
-Specifically, by adjusting a few parameters in specific Transformer layers, for the knowledge queries in $\mathcal{D}_{f}$, we alter the first token of their responses to [UNL], leveraging its properties to ensure the model forgets these knowledge.
-In a parallel vein, we maintain the knowledge associations in $\mathcal{D}_{r}$ to keep the LLM’s general ability.
-We conduct extensive experiments to validate the effectiveness of our UniErase.
-In fact, we consider fictitious and real-world knowledge in batch, sequential and precise unlearning scenarios for the Llama series of LLMs.
-Evaluating via multi-dimensional metrics on TOFU benchmark, UniErase significantly outperforms the previous SOTA forgetting baseline, achieving a 4.01$\sim$ 4.01$\times$ 4.01 × improvement in model ability while demonstrating better unlearning.
-Additionally, compared with prior retaining SOTA, UniErase maintains superior model ability and is 35.96% higher in forgetting efficacy.
+While the Large Language Models (LLMs) community (Guo et al., 2025; Chang et al., 2024; Wang et al., 2025a) has made significant advances in “learning” general abilities and domain-specific knowledge via pretraining and post-training (Kumar et al., 2025; Tie et al., 2025).
+Meanwhile, an equally crucial research direction is the complementary concept of LLM “unlearning” (Liu et al., 2025; Geng et al., 2025), which serves to address critical issues related to hallucination (Huang et al., 2025), privacy (Yan et al., 2024), and safety (Wang et al., 2025a)—including updating outdated knowledge, removing private information, and eliminating harmful contents (Lu et al., 2024; Zhang et al., 2024c; Xu, 2024).
+The core objectives of ideal unlearning is to enable LLMs, trained on trillion-token corpora, to only forget a specific data subset (the forgetting set) without compromising their general knowledge (the retaining set) and capabilities (Si et al., 2023; Maini et al., 2024).
+Given the prohibitive computational cost of retraining LLMs from scratch while excluding the forgetting set, fine-tuning (FT) techniques has emerged as the predominant unlearning implementation (Maini et al., 2024; Yuan et al., 2024).
+Concretely, FT-based unlearning can be broadly categorized into two paradigms: (I) Targeted unlearning deliberately modifies LLMs’ outputs of the forgetting set in controlled and specified manners, such as “I don’t know”-like expressions (Wei et al., 2021; Rafailov et al., 2023); (II) Untargeted unlearning shifts the responses away from the original outputs but without specifying a particular direction, like irrelevant answers (Maini et al., 2024; Zhang et al., 2024b).
+These two paradigms both employ carefully designed loss functions with distinct objectives for the forgetting set (forgetting loss) and retaining set (retaining loss), respectively, thereby enabling knowledge erasure and retention (Yuan et al., 2024; Wang et al., 2025c; b).
+However, fine-tuning inherently requires sufficient data volume to achieve effective optimization without overfitting, and the forgetting loss and retaining loss present competing objectives (Yuan et al., 2024; Geng et al., 2025).
+Besides, current FT-based unlearning, due to cost considerations, limits the retaining set to distributions near the forgetting set, which cannot represent the broad LLM knowledge (Maini et al., 2024).
+Consequently, these render two critical goals fundamentally challenging: ❶ precise unlearning for a small or even single-entry forgetting set, or ❷ balanced unlearning that concurrently preserves general abilities and knowledge while ensuring high unlearning efficacy on the forgetting set (Veldanda et al., 2024; Qu et al., 2024).
+Our empirical experiments across 8 baselines validate the second dilemma: for batch unlearning, the best-forgetting method loses 80.28% of the general abilities, while the best-retaining baseline only forgets $\sim$ half of the target data.
+Figure 1: Our UniErase achieves the most balanced unlearning performances (Left) and maintains consistently high general capabilities (Middle), delivering rapid processing and high precision (Right).
+In this paper, we aim to tackle these two critical issues for LLM knowledge unlearning, especially the balance challenge.
+To this end, we propose UniErase, a novel unlearning paradigm that balances unlearning efficacy and model abilities with dual-high performances, while supporting effective and efficient precise unlearning.
+Technically, UniErase consists of two innovative techniques: the Unlearning Token and the Unlearning Edit (Udit).
+We first introduce the unlearning token that concretizes the concept of forgetting into a tangible entity and points to a representational space that encodes unlearning semantics.
+Specifically, the unlearning token directs the autoregressive prediction process to generate predefined forgetting responses for any input sequence that terminates with this token.
+To obtain it without affecting any other generation, we create and optimize a new meta token (Li & Liang, 2021; Lester et al., 2021) only in the embedding space of the LLM, with other parameters frozen.
+Building upon this, we further propose Udit, a data-volume-independent method (therefore supporting precise unlearning) that directly modifies model parameters to establish associations between the forgetting set and unlearning token, thus realizing unlearning via its directing property.
+More importantly, Udit employs the null space projection technique (Fang et al., 2024) to ensure parameter updates remaining orthogonal to the LLMs’ existing knowledge representations, effectively preserving the retaining set and even general capabilities.
+In contrast to FT-based unlearning, UniErase pioneers the modeling of LLM unlearning as a knowledge editing problem.
+We solve the problem that current LLM editing frameworks only support entity concept editing (Wang et al., 2024; Zhang et al., 2024a) via the unlearning token, and further propose Udit to truly achieve precise and balanced LLM unlearning.
+To validate the effectiveness of UniErase, following previous works (Yuan et al., 2024; Zhang et al., 2024b), we conduct extensive experiments on different scales of the Llama-3 (Dubey et al., 2024) LLMs.
+Actually, with 8 baselines, we consider both fictitious and real-world knowledge in batch, sequential and precise unlearning scenarios (as illustrated in Figure 1).
+Evaluating via multi-dimensional metrics on the TOFU (Maini et al., 2024) benchmark, UniErase significantly outperforms the previous best-forgetting baseline, attaining $4.01\times$ performances in maintaining general knowledge and abilities while demonstrating better unlearning.
+Additionally, compared with the best-retaining baseline, UniErase preserves superior LLM abilities and is 35.96% higher in unlearning efficacy.
 In summary, our contributions can be listed as follows:
-  • Brand-new Paradigms.
-Building upon UniErase, we propose a novel paradigm that achieves efficient unlearning by directly modifying model parameters instead of multi-round fine-tuning, significantly expanding the scope of future research in the current unlearning technology stack.
-  • Balanced Trade-off. Compared to previous SOTA methods, UniErase achieves more thorough unlearning and better general ability preserving, boosting the practical usability of LLM unlearning.
-  • Generalized Scenarios. UniErase achieves the best performances across batch, sequential and precise unlearning for fictitious and real-world knowledge, covering nearly all unlearning cases.
+  • Brand-new Paradigm.
+Our proposed UniErase represents a novel unlearning paradigm that exhibits outstanding performances by directly modifying LLM parameters instead of multi-round fine-tuning, significantly expanding the scopes of future research in the unlearning community.
+  • Dual-high Balance. UniErase achieves more thorough unlearning with better retention for general knowledge and abilities, boosting the practical usability of LLM unlearning.
+  • Generalized Scenarios. UniErase performs superbly across batch, sequential and especially precise unlearning for fictitious and real-world knowledge, covering diverse unlearning tasks.
 2 Related Works
 Machine Unlearning.
-The concept of machine unlearning from traditional models bourtoule2021machine ; chen2022graph ; nguyen2022survey is emerging as a rising approach for LLM alignment research liu2025rethinking ; thaker2024position .
-Its primary goal is to enable LLMs to forget a subset $\mathcal{D}_{f}$ of training data $\mathcal{D}$ (e.g., privacy or harmful knowledge) and maintain the abilities on a retain set $\mathcal{D}_{r}\subset\mathcal{D}$, without the high cost of retraining geng2025comprehensive .
-Current mainstream approaches relying on fine-tuning and designing various loss functions for different objectives (forget $\mathcal{D}_{f}$ and retaining $\mathcal{D}_{r}$).
-Specifically, GD liu2022continual reduces the probability of generating responses to $\mathcal{D}_{f}$ by ascending gradients, and introduces another loss to constrain the deviation between the unlearned model and the original one liu2022continual .
-Meanwhile, NPO zhang2024negative , based on preference optimization rafailov2023direct , achieves unlearning by solely using $\mathcal{D}_{f}$ as negative preference, ignoring the positive term.
-Other works, such as RMU huu2024effects and LUNAR shen2025lunar , employ steering-vector-like approaches cao2024personalized to forcibly modify hidden states (but still via fine-tuning) and redirect $\mathcal{D}_{f}$ toward inability space.
-Additionally, SPULbhaila2024soft makes preliminary attempts in unlearning by adding soft prompts li2021prefix ; lester2021power during inference to manipulate model responses, but without modifying parameters to achieve “essential forgetting”.
-Model Editing.
-LLMs may contain outdated or even incorrect information huang2025survey ; tonmoy2024comprehensive .
-However, like unlearning, retraining to remove is costly, while fine-tuning overfits with limited updates.
-Thus, the light and precise model editing wang2024knowledge ; he2024llms is proposed for truthfulness huang2024can , and safety li2024precision ; chen2024can ; li2024badedit .
-Specifically, interpretability study treats LLMs’ two-layer Multilayer Perceptrons (MLP) as lookup tables for storing knowledge geva2020transformer ; zhao2021non .
-So Early methods like ROME meng2022locating and MEMIT meng2022mass introduce the locate-then-edit paradigm by solving closed-form optimization on the MLP’s down-projection matrix.
-AlphaEdit fang2024alphaedit further efficiently preserves other knowledge by projecting updates into the null space of the original knowledge representation.
-AnyEdit jiang2025anyedit decomposes edits into sub-operations for more complex knowledge.
-However, editing itself cannot accomplish unlearning directly: it requires knowledge represented in (subject, relation, object) triples and updates the object to a new value, but there exists no object token that corresponds to the abstract concept of unlearning.
-Our UniErase introduces unlearning-specific editing that integrates with existing methods.
+The concept of machine unlearning (Bourtoule et al., 2021) from traditional models (Chen et al., 2022; Nguyen et al., 2022) is emerging as a rising research topic for LLMs (Liu et al., 2025; Thaker et al., 2024).
+Its primary goal is to enable LLMs to forget a subset $\mathcal{D}_{f}$ (e.g., privacy or harmful knowledge) of the training data $\mathcal{D}$ and maintain the knowledge on a retaining set $\mathcal{D}_{r}\subset\mathcal{D}$, without the high cost of retraining (Geng et al., 2025).
+Mainstream approaches relying on the fine-tuning techniques and designing various loss functions for different objectives to simultaneously forget $\mathcal{D}_{f}$ and retain $\mathcal{D}_{r}$.
+For example, GD (Liu et al., 2022) reduces the probability of generating outputs in $\mathcal{D}_{f}$ by ascending gradients, and introduces another loss to constrain the deviation.
+Meanwhile, NPO (Zhang et al., 2024b), inspired by preference optimization (Rafailov et al., 2023), realizes unlearning by solely using $\mathcal{D}_{f}$ as negative preferences, ignoring the positive terms.
+Other works, such as RMU (Huu-Tien et al., 2024) and LUNAR (Shen et al., 2025), employ steering-vector-like approaches (Cao et al., 2024) to forcibly modify hidden states and redirect $\mathcal{D}_{f}$ toward the inability space.
+Additionally, SPUL (Bhaila et al., 2024) makes preliminary attempts in unlearning by adding soft prompts (Li & Liang, 2021; Lester et al., 2021) during inference to manipulate model responses, but without modifying parameters to achieve essential forgetting.
+Model Edit.
+LLMs may contain outdated, incorrect or even harmful information (Huang et al., 2025; Tonmoy et al., 2024).
+However, similar to unlearning, retraining for knowledge updates is costly, while fine-tuning overfits for precise scenarios.
+Thus, the model edit techniques (Wang et al., 2024; He et al., 2024) are proposed for truthfulness (Huang et al., 2024), and safety (Chen et al., 2024; Li et al., 2024).
+Early methods like ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b) introduce the locate-then-edit paradigm by modifying the down-projection matrices in the LLMs’ Multi-layer Perceptron (MLP) module.
+AlphaEdit (Fang et al., 2024) further preserves other knowledge via the null space projection operation.
+However, recent unlearning surveys like  (Liu et al., 2025) have highlighted challenges including undefined edit objectives if directly applying editing for unlearning.
+In fact, editing itself targets at knowledge represented in the (subject, relation, object) triple formats and modifies the object to a new value (Meng et al., 2022a; Wang et al., 2024; Li et al., 2025), yet no single object token exists for the abstract concept of unlearning.
+Our UniErase addresses these fundamental issues by introducing Udit with the unlearning token.
 3 Preliminaries
 Notations.
-We consider an LLM as a probabilistic language model with parameters $\theta$, denoted as $\pi_{\theta}$.
-The target set for forgetting and the set of retained knowledge are represented as $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, where typical elements of both are question and answer pairs in the form of $d=(q,a)$.
+We refer to an LLM with parameters $\theta$ as $\pi_{\theta}$.
+The target knowledge for the forgetting set and retaining set are represented as $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, where typical elements of both are question $q$ and answer $a$ pairs in the form of $d=(q,a)$.
 In addition, we denote the set of real numbers as $\mathbb{R}$, and the set of real number tensors with dimensions $(d_{1},...,d_{n})$ as $\mathbb{R}^{d_{1}\times...\times d_{n}}$.
-The symbol $\oplus$ represents the concatenation operation of token sequences.
 Unlearning Target.
-For an LLM $\pi_{\theta}$ trained with dataset $\mathcal{D}$, unlearning process aims to make the model forget the contents in $\mathcal{D}_{f}\subset\mathcal{D}$, such that it behaves as if it were trained solely on $\mathcal{D}\setminus\mathcal{D}_{f}$.
-In a parallel vein, unlearning must preserve the model’s knowledge on other datasets and its general capabilities.
-Similar to the trade-off between safety and helpfulness in LLM optimization varshney2023art , unlearning involves a balance between unlearning efficacy and model ability, formulated as:
+For an LLM $\pi_{\theta}$ trained with dataset $\mathcal{D}$, unlearning aims to make the model forget the contents in $\mathcal{D}_{f}$ as if it were trained solely on $\mathcal{D}\setminus\mathcal{D}_{f}$.
+In a parallel vein, unlearning must preserve the model’s knowledge in $\mathcal{D}_{r}$ and even broader knowledge with general capabilities.
+Similar to the trade-off between harmless and helpfulness in LLM safety alignment (Varshney et al., 2023), unlearning involves a balance between the unlearning efficacy and model ability, formulated as:
 $\pi_{\theta}^{*}=\arg\max_{\pi_{\theta}}\mathbb{E}\big[\sum_{d\in\mathcal{D}_{f}}\mathrm{Forget}(d;\pi_{\theta})+\sum_{d\in\mathcal{D}_{r}}\mathrm{Ability}(d;\pi_{\theta})\big],$ (1)
-where “$\mathrm{Forget}$” and “$\mathrm{Ability}$” are the standards or metrics for unlearning efficacy and model ability.
-Mainstream Unlearning Strategies.
-To achieve the goal in Eq 1, mainstream unlearning methods design distinct forgetting/retaining losses $l_{f}/l_{r}$ on $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, respectively, sometimes using the original model $\pi_{\theta}^{\text{ref}}$ as a reference.
-We unify their loss designs as follows, with $\beta$ and $\gamma$ as the weights:
+where $\mathrm{Forget}$” and $\mathrm{Ability}$” are the standards or metrics for unlearning efficacy and model ability.
+Mainstream Unlearning Paradigms.
+To achieve the goal in Eq.
+1, current FT-based unlearning design diverse forgetting losses $l_{f}$ and retaining losses $l_{r}$, respectively, sometimes using the original model $\pi_{\theta}^{\text{ref}}$ as a reference.
+We unify their loss designs as follows, with $\beta$ and $\gamma$ as trade-off weights:
 $\arg\min_{\pi_{\theta}}=\beta\underbrace{\mathbb{E}_{(q,a)\sim\mathcal{D}_{f}}\big[l_{f}(q\mid a;\pi_{\theta},\pi_{\theta}^{\text{ref}})\big]}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{forgetting term}}}+\gamma\underbrace{\mathbb{E}_{(q,a)\sim\mathcal{D}_{r}}\big[l_{r}(q\mid a;\pi_{\theta},\pi_{\theta}^{\text{ref}})\big]}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{retaining term}}}.$ (2)
-In Eq 2, the forgetting term is designed to make the model forget the content on $\mathcal{D}_{f}$, while the retaining term aims to preserve the knowledge on $\mathcal{D}_{r}$.
-However, current methods typically select the neighboring knowledge of $\mathcal{D}_{f}$ as $\mathcal{D}_{r}$, and can not encompass diverse general knowledge due to high time cost.
+In Eq.
+2, the forgetting term is designed to make the model forget the contents on $\mathcal{D}_{f}$, while the retaining term aims to preserve the knowledge on $\mathcal{D}_{r}$.
+Current methods typically select $\mathcal{D}_{r}$ to be the neighboring knowledge of $\mathcal{D}_{f}$, which can not encompass diverse general knowledge and abilities.
 In Appendix B, we introduce the specific forms of various $l_{f}$ and $l_{r}$ in detail.
-4 Methodology
-In this section, we initially employ mathematical logic to expound upon the fundamental principles of UniErase ($\triangleright$ Sec 4.1), as demonstrated in Figure 2.
-Then, we elaborate on the methodology of training an unlearning token via incorporating a minimal number of parameters ($\triangleright$ Sec 4.2).
-Additionally, we discuss the process of precisely modifying parameters for the forgetting targets ($\triangleright$ Sec 4.3), aiming at achieving an optimal equilibrium between high unlearning efficacy and low ability loss ($\triangleright$ Sec 4.4).
+4 Perform Unlearning Edit with Unleanring Token
+In this section, we first introduce the Unlearning Logical Chain to expound upon the fundamental principles of UniErase ($\triangleright$ Section 4.1), as demonstrated in Figure 2.
+Then, we propose the unlearning token and elaborate on the techniques to obtain it via incorporating a minimal number of parameters ($\triangleright$ Section 4.2).
+Subsequently, we introduce Udit to modify parameters for the unlearning targets ($\triangleright$ Section 4.3), achieving balanced unlearning performances while supporting precise unlearning.
 4.1 Unlearning Logical Chain
 Given an LLM $\pi_{\theta}$, for an input token sequence $q=[x_{1}x_{2}...x_{n}]$, we assume the output token sequence is $a=[y_{1}y_{2}...y_{m}]$.
-Then we abstract this generation process as mathematical logic derivations:
+Then we abstract this generation process as a mathematical logic derivation:
 $x_{1}x_{2}...x_{n}\xRightarrow{\pi_{\theta}}y_{1}\xRightarrow{\pi_{\theta}}y_{2}\xRightarrow{\pi_{\theta}}...\xRightarrow{\pi_{\theta}}y_{m},$ (3)
@@
 Proposition 1.
-The Unlearning Token is a novel token trained on $\mathcal{D}_{f}$, designed to constrain $\pi_{\theta}$ such that its subsequent token generation redirects targeted concept forgetting or induced unawareness.
-Specifically, for $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$, the unlearning token [UNL] is defined as follows:
-$x_{1}x_{2}...x_{n}\oplus\mathrm{[UNL]}\xRightarrow{\pi_{\theta}}y_{\text{idk}},\quad y_{\text{idk}}\in\mathcal{D}_{\text{idk}},$ (4)
-where $\mathcal{D}_{\text{idk}}$ is the set of token sequences that express forgetting or ignorance (e.g., “I do not know”).
-Building upon Proposition 1, we have learned the [UNL] token.
-However, when only $q$ is provided as input, the model still generates original normal response rather than $y_{\text{idk}}$.
-Therefore, we need to modify model parameters so that for any $q\in\mathcal{D}_{f}$, the first token of the generated response is [UNL], thereby internalizing “forgetting $q$ with [UNL]” as part of the model’s knowledge or “cognition”.
-Model editing, as an evolving approach to edit model knowledge without ability degradation fang2024alphaedit ; meng2022locating , basically meets our unlearning requirements if adapted properly.
+The Unlearning Token (denoted as [UNL]) is a novel token, designed to direct the LLM’s subsequent token generation to specific forgetting expressions.
+We refer to the token concatenation operator as $\oplus$.
+Then, for any $(q,a)\in\mathcal{D}$, [UNL] satisfies the following property:
+$x_{1}x_{2}...x_{n}\oplus\mathrm{[UNL]}\xRightarrow{\pi_{\theta}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}\quad\land\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{\theta}}a,$ (4)
+where operation $a\land b$ means that both $a$ and $b$ should be satisfied and $\mathcal{D}_{\text{idk}}$ contains different token sequences that express the semantics of forgetting or ignorance.
+Specifically, Eq.
+4 stipulates that the newly acquired unlearning token should exclusively direct the model toward the forgetting semantic space when employed as a suffix, while preserving normal knowledge retrieval capabilities otherwise.
+In Proposition 1, we have defined the [UNL] meta token.
+However, when only $q$ is provided as input, the LLM still generates original normal responses rather than $y_{\text{idk}}$.
+To realize unlearning, we need to modify model parameters so that: for any $q$, its next token prediction is [UNL], thereby internalizing “forgetting $q$ with [UNL]” as the LLM’s inherent knowledge.
 To this end, we propose:
 Proposition 2.
-Unlearning Editing utilizes model editing to modify only a small set of parameters ($|\Delta\theta|\ll|\theta|$), enabling the model to forget specified intrinsic knowledge.
-For $q=[x_{1}x_{2}...x_{n}]\in\mathcal{D}_{f}$ with the obtained [UNL], the parameter $\theta$ is updated with $\Delta\theta$, formally defined as:
-$\theta\leftarrow\theta+\Delta\theta,\quad|\Delta\theta|\ll|\theta|,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}.$ (5)
-Derivation: Eq 4 ensures that [UNL] can direct the input to the unlearning space (generating responses in $\mathcal{D}_{\text{idk}}$) when used as a suffix, while Eq 5 changes the next token of the normal response of the input sequence to [UNL].
-Based on these two propositions, we can derive the following Unlearning Logical Chain to adjust model parameters for achieving our UniErase unlearning:
-$\theta\leftarrow\theta+\Delta\theta,\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\xRightarrow{\mathrm{\pi_{{\theta}}}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}.$ (6)
-This chain enables us to redirect $\mathcal{D}_{f}$ into the unlearning space and leverage model editing to preserve general capabilities. However, the following challenges arise for implementation:
-Challenge I: How to teach LLMs to learn and generate [UNL] without overfitting?
-Since the off-the-shelf soft prompt li2021prefix ; lester2021power ; bhaila2024soft optimization only considers meta tokens as input prefix or suffix without truly introducing a new token.
-To represent the unlearning space, we need to teach the model the meaning of the unlearning token [UNL].
-Besides, editing in Eq 5 requires LLM’s ability to generate the target token (as one step involves maximizing its generation probability).
-Therefore, we should optimize the model to generate [UNL].
-However, the samll $|\mathcal{D}_{f}|$ may lead to overfitting, where the LLM incorrectly holds $q\xRightarrow{\pi_{\theta}}\mathrm{[UNL]}$ for $(q,a)\notin\mathcal{D}_{f}$ (even for $(q,a)\in\mathcal{D}_{r}$) before editing.
-Challenge II: How to enhance the robustness of [UNL]?
-The above derivation is based on model’s continuing generation, but aligned LLMs in practical applications typically append chat templates to input (e.g., user/assistant role, demonstration in Appendix C), which may cause [UNL] fail to induce $\mathcal{D}_{\text{idk}}$.
-On the other hand, in Eq 5, even if the scale of parameter change is small ($|\Delta\theta|\ll|\theta|$), $\Delta\theta$ may turn Eq 4 not true for the updated model.
-Thus, for the derivation in Eq 6, we need to ensure that [UNL] remains effective after parameter modifications, successfully serving as a token to induce $y_{\text{idk}}$.
-Figure 2: Overview framework of UniErase. We demonstrate current unlearning methods and our UniErase, consisting of unlearning logical chain, unlearning token and unlearning editing.
-4.2 Generating Unlearning Token (Proposition 1)
-In this section, with Figure 2, we first introduce how to learn the unlearning token [UNL] by introducing a single vector both in the embedding and unembedding matrix of the LLM, and then propose specific techniques to address the two main challenges discussed in the above analysis.
-4.2.1 Optimization Target for [UNL] (☞ Challenge I)
-Given the embedding matrix of $\pi_{\theta}$ as $E\subset\theta$ and $E\in\mathbb{R}^{n\times d}$, where $n$ is the vocabulary size and $d$ is the embedding dimension, our goal is to add a new token [UNL] and optimize its corresponding embedding vector $e_{u}\in\mathbb{R}^{d}$ to link it with the forgetting space in the LLM via the follow loss:
-$\arg\min_{e_{u}}\mathcal{L}_{ul}=\arg\min_{e_{u}}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})\big{]},$ (7)
+Unlearning Editing (Udit) modifies only a small set of parameters $\Delta\theta$, enabling the LLM to forget specified knowledge.
+For $\forall(q,a)\in\mathcal{D}_{f}$ and $\forall(q^{\prime},a^{\prime})\in\mathcal{D}\setminus\mathcal{D}_{f}$, Udit ensures that:
+$\left(|\Delta\theta|\ll|\theta|,\quad\theta\leftarrow\theta+\Delta\theta\quad\text{s.t.}\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\right)\quad\land\quad\left(q^{\prime}\xRightarrow{\pi_{{\theta}}}a^{\prime}\right)$ (5)
+According to Eq.
+5, Udit must demonstrate the ability to alter the subsequent token prediction of target unlearning contents to [UNL] via sparse parameter updates, while maintaining intact knowledge retrieval and response capabilities for non-target contents.
+Derivation: Grounded in the aforementioned propositions, we establish the following Unlearning Logical Chain, which directly modifies LLM parameters to accomplish efficient targeted unlearning without compromising the model’s retained knowledge and general capabilities:
+$\left(q^{\prime}\xRightarrow{\pi_{{\theta}}}a^{\prime}\right)\quad\land\quad\left(\theta\leftarrow\theta+\Delta\theta\quad\text{s.t.}\quad x_{1}x_{2}...x_{n}\xRightarrow{\pi_{{\theta}}}\mathrm{[UNL]}\xRightarrow{\mathrm{\pi_{{\theta}}}}y_{\text{idk}}\in\mathcal{D}_{\text{idk}}\right)$ (6)
+This chain demonstrates the core spirits of UniErase, enabling us to realize unlearning on $\mathcal{D}_{f}$ via directing $a\in\mathcal{D}_{F}$ to $y_{\text{idk}}$, while preserving other untargeted generation $q^{\prime}\to a^{\prime}$.
+Figure 2: Paradigm of UniErase and comparison with mainstream FT-based unlearning.
+4.2 Unlearning Token
+In this section, we present the specific techniques for deriving the unlearning token that fulfill the requirements in the Unlearning Logical Chain.
+In essence, the special token must satisfy three key criteria: redirecting arbitrary knowledge toward the forgetting space (Eq.
+4), maintaining the normal response generation for other knowledge domains (Eq.
+4), and being a generatable token (Eq.
+5) rather than appearing only at the input end like prefix tuning (Li & Liang, 2021; Bhaila et al., 2024).
+4.2.1 Optimization to Attain Unlearning Token
+Let $E\in\mathbb{R}^{n\times d}\subset\theta$ and $U\in\mathbb{R}^{n\times d}$ denote the embedding and unembedding matrices of the LLM $\pi_{\theta}$, respectively, where $n$ represents the vocabulary size and $d$ denotes the model dimension.
+We expand both $E$ and $U$ by incorporating two additional row vectors: $e_{0}\in\mathbb{R}^{d}$ and $u_{0}\in\mathbb{R}^{d}$ ($E\leftarrow E\cup e_{0}$ and $U\leftarrow U\cup u_{0}$), which correspond to the encoding and decoding representations for the unlearning token [UNL], respectively.
+Subsequently, we optimize the following objective to learn [UNL]:
+$\arg\min_{(e_{u},\hat{e}_{u})}\mathbb{E}_{(q,a)\sim\mathcal{D}_{f},a^{\prime}\sim\mathcal{D}_{\text{idk}}}\big[-\log P(a^{\prime}\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta}^{u})-\log P(a\mid q;\pi_{\theta}^{u})\big],$ (7)
 where $P(x\mid y;\pi_{\theta})$ means the conditional probability of $\pi_{\theta}$ generating output $x$ when giving input $y$.
-Eq 7 ensures that upon encountering [UNL], the model directs the response $a$ (of input $q$) to the forgetting space, with $E\leftarrow E\cup e_{u}\in\mathbb{R}^{(n+1)\times d}$.
-Prompt Tuning lester2021power shares similar concepts with Eq 7, but to enable the model to generate [UNL], we make some refinements and actually consider:
-$\arg\min_{(e_{u},\hat{e}_{u})}\mathbb{E}_{q\sim\mathcal{D}_{f},a\sim\mathcal{D}_{\text{idk}}}\big{[}-\log p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta}^{u})\big{]},$ (8)
-where $\hat{e}_{u}$ denotes a newly added unembedding vector for [UNL] in the unembedding matrix $U\in\mathbb{R}^{n\times d}$ ($U\leftarrow U\cup\hat{e}_{u}\in\mathbb{R}^{(n+1)\times d}$), and $\pi_{\theta}^{u}$ is the model initializing $e_{u}$ and $\hat{e}_{u}$ randomly.
-Unlike Eq 7, which only expands $E$, Eq 8 jointly expands both $E$ and $U$.
-This supervises the LLM to output unlearning expressions $\mathcal{D}_{\text{idk}}$ after [UNL] and equips the model with the ability to generate [UNL] for subsequent edits.
-However, Eq 8 risks overfitting, as detailed in Chanllenge I in section 4.1.
-To address this, we adopt an empirically validated technique: sharing the parameters of [UNL] added to $E$ and $U$ (by setting $e_{u}=\hat{e}_{u}$).
-This approach has also been employed in the pre-training of some LLMs touvron2023llama .
-We provide experiments and analysis on solving the overfitting problem in Appendix D.
-4.2.2 Robustness Enhancement of [UNL] (☞ Challenge II)
-Template Robustness refers to the property where [UNL] can still steer the LLM response toward $\mathcal{D_{\text{ign}}}$ even when additional prefixes/suffixes (such as chat template function $\mathrm{Template}$) are prepended/appended to the original input $q$.
-Denoting $p_{0}=p(a\mid q\oplus\mathrm{[UNL]};\pi_{\theta})$, we formulate it as:
-$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}|\mathrm{Template}(q\oplus\mathrm{[UNL])};\pi_{\theta})|\big{]}.$ (9)
-Parameter Robustness characterizes the feature that after subsequent model edits (as described in Eq 5), [UNL] can still steer the input toward $a\in\mathcal{D_{\text{ign}}}$.
-Supposing the distribution of parameter perturbation $\epsilon$ is $\mathcal{N}$, the objective can be formalized as follows:
-$\arg\min_{\pi_{\theta}}\mathbb{E}_{q\sim\mathcal{D}_{f},\epsilon\sim\mathcal{N}}\big{[}|p_{0}-p(a\oplus\mathrm{[UNL]}\mid q\oplus\mathrm{[UNL]};\pi_{\theta+\epsilon})|\big{]}.$ (10)
-Directly solving Eq 9 or 10 is challenging.
-Therefore, we adopt a reverse approach to enhancing robustness through extra optimization on [UNL] based on Eq 8.
-Specifically, after initial optimization, we mix $\mathcal{D}_{f}$ and $\mathrm{Template}(\mathcal{D}_{f})$ and optimize $(e_{u},\hat{e}_{u})$ again for Eq 9.
-Then, since editing only modifies the down-projection matrix $W_{\text{dp}}^{i}$ of the $i$-th MLP layer, for Eq 10, we further optimize [UNL] via adversarial perturbation: $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\alpha f(W_{\text{dp}}^{i})$, where $f$ is a distribution function.
-4.3 Performing Unlearning Edits (Proposition 2)
-Upon obtaining [UNL], we then edit the MLP layer of the LLM so that for $\forall(q,a)\in\mathcal{D}_{f}$, the mapping $q\xRightarrow{\pi_{{\theta}}}a$ is modified to $q\xRightarrow{\pi_{{\theta}}}a^{*}\in\mathcal{D}_{\text{idk}}$ for the internalization of unlearning targets.
+Eq.
+7 ensures that upon encountering [UNL], the model directs the original response $a$ (of input $q$) to the forgetting space, yielding $a^{\prime}\in\mathcal{D}_{\text{idk}}$.
+Concurrently, the LLM is also forced to learn to generate $a^{\prime}\oplus\mathrm{[UNL]}$, satisfying the requirement specified in Eq.
+5 that the LLM possesses the capability to generate [UNL].
+Furthermore, since we only introduce additional parameters in $E$ and $U$, while keeping all other parameters frozen, the aforementioned [UNL] optimization does not interfere with normal response generation for other knowledge domains.
+4.2.2 Robustness Enhancement of Unlearning Token
+Parameter Robustness.
+In the Unlearning Logical Chain, $\Delta$ in Udit may render the previously learned unlearning token ineffective, so we need to enhance its robustness against slight parameter perturbations.
+While directly incorporating constraints into Eq.
+7 to achieve this is challenging, we leverage the fact that Udit’s target parameters are confined to the down projection matrices $W_{dp}$ within the MLP module.
+Therefore, in the optimization process of Eq.
+7, we introduce the following perturbations in the LLM to improve the parameter robustness of the resulting unlearning tokens:
+$W_{\text{dp}}^{i}\in\theta\leftarrow W_{\text{dp}}^{i}+\alpha f(W_{\text{dp}}^{i})\cdot W_{\text{dp}}^{i},$ (8)
+where parameter $\alpha$ controls the intensity and $f$ is a function mapping $W_{\text{dp}}^{i}$ to a scalar.
+4.3 Unlearning Edit
+With [UNL] obtained, we propose Udit to bridge it with the knowledge to be unlearned, ensuring the internalization of unlearning targets.
+Following previous model editing techniques (Meng et al., 2022a; b), we target at the down projection matrices $W_{\text{dp}}^{i}$ in the MLP module.
 As shown in Figure 2, the $i$-th MLP layer performs the following computation ($\sigma$ and $\gamma$ are activation functions):
@@
 where $h^{i}$, $a^{i}$, and $m^{i}$ represent the hidden state, attention and MLP output in the $i$-th layer, respectively.
-Unlearning editing only modifies $W_{\text{dp}}^{i}\leftarrow\tilde{W}_{\text{dp}}^{i}$ to ensure $\tilde{W}_{\text{dp}}^{i}k^{*}=v^{*}$ holds, where $k^{*}$ represents the corresponding hidden state of $q$, and $v^{*}$ maximizes the prediction probability of [UNL] as the next token.
-We provide the details on how to get these new $k^{*}$ and $v^{*}$ for each $(q,a)$ in Appendix E.
-In other words, we should then build new mappings from $W_{\text{dp}}^{i}k=v\to W_{\text{dp}}^{i}k^{*}=v^{*}$ to implement unlearning edits described in Eq 5.
-For each $(q,a)$ in $D_{f}$ and $D_{r}$, we stack their corresponding $k^{*}$, $v^{*}$ vectors into matrices $K_{f}$, $V_{f}$ and $K_{r}$, $V_{r}$, respectively.
-Following ROME meng2022locating , we build the new mappings by updating $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\Delta^{*}$ to approximate the overall unlearning standard in Eq 1 via:
-$\Delta^{*}=\arg\min_{\Delta}\big{(}\underbrace{\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|(W_{\text{dp}}^{i}+\Delta)K_{r}-V_{r}\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\}\big{)}.$ (12)
-In Eq 12, for all $(q,a)\in\mathcal{D}_{f}$, the forget term modifies the first token of the response $a$ to [UNL], while the retain term ensures that all $(q,a)\in\mathcal{D}_{r}$ retain their original input-output pairs.
-In practice, to further maintain model’s ability, we include a sample of general knowledge $\mathcal{D}_{g}$ in $\mathcal{D}_{r}$ ($|\mathcal{D}_{g}|\gg|\mathcal{D}_{f}|$), which other unlearning methods unable to involve due to high computational costs yuan2024closer ; zhang2024negative .
-Then, to solve Eq 12, we can quickly get its closed-form solution, without optimization process:
+Udit exclusively updates $W_{\text{dp}}^{i}\leftarrow\tilde{W}_{\text{dp}}^{i}$ to satisfy the new association $\tilde{W}_{\text{dp}}^{i}k^{*}=v^{*}$, where $k^{*}$ corresponds to the hidden state of targeted knowledge $q\in\mathcal{D}_{f}$, and $v^{*}$ is optimized to maximize the prediction probability of [UNL] as the next token when the input is $q$.
+In other words, Udit builds new knowledge mappings from the original $W_{\text{dp}}^{i}k=v\to W_{\text{dp}}^{i}k^{*}=v^{*}$ to ensure the next token prediction for $q\in\mathcal{D}_{f}$ is modified to [UNL] (Eq.
+5).
+The detailed procedures for obtaining these $k^{*}$ and $v^{*}$ for each knowledge-answer pair $(q,a)$ are provided in Appendix C.
+We construct the unlearning matrices by stacking the corresponding key and value vectors.
+Specifically, for each $(q,a)\in D_{f}$ to be unlearned, we stack their corresponding unlearning key vectors $k^{*}$ and value vectors $v^{*}$ into matrices ($K_{f}$, $V_{f}$), respectively.
+Similarly, for each $(q,a)\in D_{r}$ to be retained, compile their normal key and value vectors into matrices ($K_{r}$, $V_{r}$).
+Then, we propose the core technique of Udit: by updating $W_{\text{dp}}^{i}\leftarrow W_{\text{dp}}^{i}+\Delta^{*}$, we construct new mappings between $q\in\mathcal{D}_{f}$ and [UNL], while preserving the retrieval of other knowledge ($q\in\mathcal{D}_{r}$) to approximate the unlearning objective described in Eq.
+1.
+The parameter update $\Delta^{*}$ is optimized via:
+$\Delta^{*}=\arg\min_{\Delta}\big(\underbrace{\|(W_{\text{dp}}+\Delta)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|(W_{\text{dp}}^{i}+\Delta)K_{r}-V_{r}\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}retain term}}}\big).$ (10)
+In Eq.
+10, for all $(q,a)\in\mathcal{D}_{f}$, the forget term modifies the first token of the response $a$ to [UNL], while the retain term ensures that all $(q,a)\in\mathcal{D}_{r}$ retain their original input-output pairs.
+Through mathematical derivation (provided in Appendix C), we can quickly get its closed-form solution:
 $\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}(K_{r}K_{r}^{T}+K_{f}K_{f}^{T})^{-1}.$ (11)
-In Appendix E, we provide a rigorous mathematical derivation for Eq 32, along with the other unlearning editing methods (derived from 12) that better redirect $D_{f}$ to [UNL] or preserve $D_{r}$.
-4.4 Balancing between Unlearning Efficacy and Model Ability
-In summary, Sec 4.2 introduces methods to obtain a robust new token as [UNL], clarifying a specific unlearning space.
-Sec 4.3 then executes unlearning edits to achieve redirecting $\mathcal{D}_{f}$ to [UNL].
-Collectively, the reasons that UniErase can balance unlearning and ability lies in:
-Effective Unlearning Logic. For unlearning efficay, the unlearning logical chain strictly ensures that unlearning token and unlearning editing work in synergy to achieve the targeted forgetting.
-Constrained Parameter Modification.
-For model ability, unlearning token only adds two new vectors and addresses overfitting, while unlearning editing modifies only several matrices in a few MLP layers while preserving the LLM’s original ability by adding the general knowledge set.
+Notably, Eq.
+11 accommodates a significantly broader $\mathcal{D}_{r}$ than FT-based unlearning methods, thereby preserving a wider range of general knowledge and capabilities.
+To this end, we include a general knowledge dataset $\mathcal{D}_{g}$ within $\mathcal{D}_{r}$ (with $|\mathcal{D}_{g}|\gg|\mathcal{D}_{f}|$), which remains computationally infeasible for other unlearning approaches (Yuan et al., 2024; Zhang et al., 2024b; Maini et al., 2024).
+Null-space Projection Unlearning.
+Inspired by AlphaEdit (Fang et al., 2024), we further optimize Udit into a null space projection formulation to further reduce the impact of unlearning on general knowledge.
+Specifically, we obtain the new parameter update $\Delta P$ by right-multiplying with matrix $P$, which projects $\Delta$ onto the null space of $K_{r}$ such that $\Delta PK_{r}=0$.
+Through straightforward calculation, the retain term in Eq.
+10 degenerates to 0 (meaning no influence on the retaining set), thus we only need to optimize the forget term.
+The new optimization objective can be formulated as:
+$\Delta^{*}=\arg\min_{\Delta}\big(\underbrace{\|(W_{\text{dp}}+\Delta P)K_{f}-V_{f}\|^{2}}_{\text{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}forget term}}}+\underbrace{\|\Delta P\|^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}constrain term}}}\big),$ (12)
+where the constraint term is incorporated to limit the magnitude of parameter updates. Through a similar derivation as presented in Eq. 11, we arrive at the following closed-form solution:
+$\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}P(K_{f}K_{f}^{T}P+I)^{-1}.$ (13)
+The complete mathematical derivation of the closed-form solution presented in Eq. 13, as well as the methodology for computing $P$, is provided in Appendix C.
 5 Experiment
-In this section, we experimentally validate the universality and effectiveness of our UniErase via multiple unlearning metrics in three scenarios: (I) Batch Unlearning ($\triangleright$ Sec 5.2) refers to enabling an LLM to forget a large forgetting dataset in a single unlearning step, which is a key scenario considered in current domain.
-(II) Sequential Unlearning ($\triangleright$ Sec 5.3) performs multiple consecutive rounds of batch unlearning, testing whether the LLM collapses under repeated application.
-(III) Precise Unlearning ($\triangleright$ Sec 5.4) considers extremely small forgetting sets to test the precision of unlearning methods, serving as a practical and meaningful complement to existing scenarios.
+In this section, we experimentally validate and analyze the effectiveness of our balanced and precise UniErase in the following three scenarios: (I) Batch Unlearning ($\triangleright$ Section 5.2) refers to making an LLM forget a large forgetting dataset in a single unlearning step.
+(II) Sequential Unlearning ($\triangleright$ Section 5.3) performs multiple rounds of unlearning tasks, testing whether unlearning methods cause LLMs to collapse for consecutive scenarios.
+(III) Precise Unlearning ($\triangleright$ Section 5.4) considers extremely small (single-entry) forgetting sets to test the precision of unlearning methods.
 5.1 Overall Settings
-Datasets and Models.
-We consider two widely adopted TOFU and RETURN liu2024learning benchmarks for fictitious and real-world unlearning, respectively.
-These benchmarks both contain serval forgetting set and corresponding and neighboring retaining set.
-To test LLMs’ general ability, we further apply MMLU hendrycks2020measuring for fact answering, TriviaQA joshi2017triviaqa for context comprehension, GSM8k cobbe2021training for math reasoning, and Human-Eval chen2021evaluating for coding.
-Following previous works, we perform unlearning on two scales of Llama-3 touvron2023llama serie model (3B and 8B).
-Notably, for fictitious unlearning, we use the TOFU-injected  [^1] https://huggingface.co/open-unlearning/tofu_Llama-3.1-8B-Instruct_full version.
-Details of these datasets and models are demonstrated in Appendix F.
+Datasets & Models.
+We consider two widely adopted TOFU (Maini et al., 2024) and RETURN (Liu et al., 2024) benchmarks for fictitious and real-world knowledge unlearning, respectively.
+They both contain serval forgetting sets and corresponding and neighboring retaining sets.
+Highlighting the retention for general abilities, we employ MMLU (Hendrycks et al., 2020) for fact answering, TriviaQA (Joshi et al., 2017) for context comprehension, GSM8k (Cobbe et al., 2021) for math reasoning, and Human-Eval (Chen et al., 2021) for coding.
+Following previous works (Maini et al., 2024; Yuan et al., 2024), we perform unlearning on the Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Touvron et al., 2023).
+For fictitious knowledge unlearning, we apply the model versions [^1] https://huggingface.co/open-unlearning/tofu_Llama-3.1-8B-Instruct_full fine-tuned on TOFU.
+More details of these datasets and models are demonstrated in Appendix D.
 Metrics.
-We consider multiple metrics across unlearning efficacy and model ability.
-For unlearning efficacy, in line with prior works maini2024tofu ; zhang2024negative ; yuan2024closer , we employ ROUGE (word-level match), Probability (ground truth likelihood), Truth Ratio (correct-incorrect preference), Token Entropy (generation diversity), Similarity (semantic similarity), and Entailment Score (factual correctness) ferrandez2008te4av .
-For $\mathcal{D}_{f}$, we use the arithmetic mean of these metrics as the overall Forget Efficacy (FE) yuan2024closer .
-For model ability, we similarly apply the above metrics to $\mathcal{D}_{r}$, computing their harmonic mean as Retain Efficacy (RE) yuan2024closer .
-On the other hand, for general abilities, we calculate accuracy (Acc), “I do not know”-like response ratio (Idk), and average response token number (Len) on relevant datasets.
-Besides, Retain Average (RA) is the mean of all RE and Acc.
-We provide the details of these metrics in Appendix G.
+We consider multiple metrics to comprehensively evaluate performances on unlearning and retention.
+For unlearning efficacy, in line with prior research (Maini et al., 2024; Zhang et al., 2024b; Yuan et al., 2024), we employ ROUGE (word-level match), Probability (ground truth likelihood), Truth Ratio (correct-incorrect preference), Token Entropy (generation diversity), Similarity (semantic similarity), and Entailment Score (factual correctness) (Ferrández et al., 2008).
+To obtain an integrated indicator, we calculate the arithmetic mean of these metrics on $\mathcal{D}_{f}$ as the overall Forgetting Efficacy (FE) (Yuan et al., 2024).
+For neighboring knowledge retention on $\mathcal{D}_{r}$, we similarly apply the above metrics and compute their harmonic mean to be the Retaining Efficacy (RE) (Yuan et al., 2024).
+Besides, for general abilities, we report the accuracy (Acc), “I do not know”-like response ratio (Idk), and average response token number (Len).
+Besides, Retaining Average (RA) is the mean of RE and all Accs, as the final metric for retention.
+We provide the details of these metrics in Appendix E.
 Baselines.
-We compare our UniErase with serveral available unlearning methods (detailed in Appendix B): GA liu2022continual , DPO maini2024tofu , NPO zhang2024negative and IDK yuan2024closer , which are the forgetting losses.
-We further combine them with retaining losses: KL maini2024tofu , GD liu2022continual and ME yuan2024closer , finally forming GA+GD, GA+KL yao2024large , NPO+KL zhang2024negative , ME+GD yuan2024closer , DPO+GD maini2024tofu , DPO+KL, IDK+AP yuan2024closer .
-We follow each baseline’s parameter settings and provide them in Appendix H.
-Besides, for the unlearning editing in our UniErase, we apply the AlphaEdit-derived fang2024alphaedit version (detailed in Appendix E).
-Table 1: Forget Efficacy (FE), Retain Efficacy (RE) and General Ability of Different Baselines on TOFU benchmark for Batch Unlearning.
+To demonstrate the effectiveness of our proposed paradigm, we evaluate UniErase against 8 FT-based unlearning baselines.
+Our comparison includes four primary forgetting losses: GA (Liu et al., 2022; Yao et al., 2024), DPO (Maini et al., 2024), NPO (Zhang et al., 2024b), and IDK (Yuan et al., 2024).
+These forgetting losses are combined with various retaining loss functions, including KL (Maini et al., 2024), GD (Liu et al., 2022), and ME (Yuan et al., 2024), resulting in the following baseline configurations: GA+GD, GA+KL, NPO+GD, NPO+KL, ME+GD, DPO+GD, DPO+KL, and IDK+AP.
+The specific parameter settings of all methods are detailed in Appendix F.
+Table 1: Batch unlearning performances of different unlearning methods for the TOFU-inject Llama-3.1-8B-Instruct.
 “Base” means the original LLM before unlearning.
-In each row of the baselines, we bold the maximum value and underline the second largest one.
-“Forget” and “Retain” is the most numerous $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ in TOFU, with “Real” as its real fact test set.
-We provide more experiment results (for the 3B-scale LLM) in Appendix I, showing similar patterns.
+In each row, we bold the maximum value and underline the second largest one.
+“Forget” and “Retain” refer to the $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ datasets in TOFU, while “Real” is the real fact test dataset in TOFU.
 Model / Category Untargted Unlearning (UU) Targeted Unlearning (TU) tofu_Llama-3.1-8B-Instruct_full GA+GD GA+KL NPO+GD NPO+KL ME+GD DPO+GD DPO+KL IDK+AP UniErase Dataset Metric Base - NIPS24 - COLM24 ICLR25 COLM24 - ICLR25 (Ours) Forget FE 10.95 58.29 62.91 58.31 59.24 78.01 79.31 79.02 58.42 79.43 Retain RE 86.34 27.47 0.00 43.38 31.73 41.92 0.00 0.00 78.03 82.32 Real RE 76.44 42.75 0.00 53.88 46.75 57.63 0.00 0.00 74.73 75.18 Acc 62.75 62.18 62.66 44.30 57.69 27.85 31.34 19.73 62.18 61.89 MMLU Idk 0.00 0.00 0.00 0.00 0.00 0.00 51.07 69.80 0.00 0.00 Len 8.55 20.14 172.8 511.8 499.7 28.41 7.03 7.41 6.32 8.68 Acc 82.49 82.22 80.53 82.44 80.66 78.97 54.17 35.81 80.47 82.75 TriviaQA Idk 0.00 0.00 0.00 0.00 0.00 0.00 26.89 50.46 0.00 0.00 Len 9.53 13.77 43.24 512.0 492.0 27.44 7.87 7.85 7.96 9.53 Acc 56.10 54.27 64.02 0.07 23.78 0.00 0.00 0.00 48.78 54.27 Human-Eval Idk 0.00 0.00 0.00 0.00 0.00 0.00 72.57 85.98 0.00 0.00 Len 61.53 66.85 88.46 316.6 205.7 18.91 22.26 15.36 60.74 61.98 Acc 69.37 75.36 77.71 53.53 56.33 38.59 0.00 0.00 59.14 71.57 GSM8k Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.00 0.00 Len 99.48 147.7 189.7 511.6 468.3 97.15 8.00 8.00 72.38 100.4 Retain Average (RA) 72.25 57.38 47.49 46.27 49.49 40.83 14.25 9.26 67.22 71.33 Retain Ratio (%) 100.0 79.41 65.73 64.04 68.50 56.51 19.72 12.81 93.04 98.73 Balance = (FE+RA)/2 41.60 57.83 55.20 52.29 54.37 59.42 46.78 44.14 62.82 75.38
 5.2 Batch Unlearning
-We conduct fundamental batch unlearning on TOFU and RETURN dataset, forgetting 400 fictitious and real-world knowledge entries at once, with results presented in Table 1 and Figure 3, respectively.
+We perform batch unlearning on the TOFU and RETURN forgetting datasets, eliminating 400 fictitious and real-world knowledge entries in a single batch operation.
+The experimental results are presented in Table 1 and Figure 3.
+We provide supporting results of another LLM in Appendix G.
+[…]
 Obs.
-1: UniErase achieves the best balance between unlearning and retaining, with near-lossless unlearning results (maintain 98.73% of LLMs’ ability).
-In Table 1, UniErase achieves the highest forget efficacy of 79.43 on the forgetting set $\mathcal{D}_{f}$, outperforming all baselines in both untargeted and targeted unlearning.
-Meanwhile, it also attains the retain efficacy of 82.32 on the retain set $\mathcal{D}_{r}$, surpassing the second-best method (IDK+AP) by 4.29, while UniErase’s unlearning efficacy is significantly higher by 35.96%.
-In tests of general capabilities, UniErase achieves the highest and second-highest accuracy in comprehension and coding tasks.
-For MMLU reasoning, it incurs only a 1.37% performance drop, matching the SOTA baselines (GA+KL, IDK+AP).
-Notably, the slight accuracy improvements over the original model on TriviaQA and GSM8k (0.26 and 2.2 $\uparrow$) align with AlphaEdit’s results fang2024alphaedit .
-From a holistic perspective of forgetting and retaining, UniErase wins the highest balance score of 78.38–1.15$\times$ and 1.71$\times$ higher than the second-best and the worst.
-Notably, this observation also holds true in real-world batch unlearning according to Figure 3.
-Obs.
-2: UniErase is entirely free from the over-unlearning problem.
-Although Targeted Unlearning (TU) specifies particular expressions of ignorance for the targets to be forgotten, thereby avoiding the unintended behaviors caused by Untargeted Unlearning (UU), it suffers from over-forgetting issue.
-Specifically, in Table 1, all UU baselines do not produce “I do not know”-like responses ($\mathcal{D}_{\text{idk}}$) on the four general ability datasets, with all $\text{Idk}=0$.
-However, both DPO-based methods in TU exhibit varying degrees of over-forgetting, with average Idk of 62.63 and 76.56, respectively.
-The most severe case occurs on GSM8k, where these two baselines respond answers in $\mathcal{D}_{\text{idk}}$ for 100% of cases.
-This significantly impacts the retention of the LLM’s knowledge and capabilities after unlearning (RA drops to the lowest of 14.25 and 9.26).
-In contrast, UniErase—the best-retaining TU baselines completely avoid the over-unlearning problem ($\text{Idk}=0$), with UniErase even achieving best unlearning efficacy (79.43).
-Ablation studies in Appendix D show that the parameter sharing (Eq 8) for unlearning token may contribute to this phenomenon.
-Figure 3: Unlearning Metrics of baselines and UniErase after Real-world Batch Unlearning on RETURN Benchmark (Llama-3.1-8B).
-Obs.
-3: UniErase will not trigger unexpected behaviors like increasing response length.
-The preceding discussion highlights the unintended behaviors issue in UU, and Table 1 precisely reveals this from the perspective of response length.
-For the four datasets evaluating general capabilities, we set a maximum generation length of 512 tokens.
-Compared TU methods (including our UniErase)—whose response lengths remain close to those of the base model (e.g., average token counts on MMLU range between 6.32 and 8.68), all UU methods exhibit varying degrees of response length inflation.
-The most severe cases are the two NPO-based methods, where NPO+GD, according to the Len metric, generates responses up to 50 $\times$ longer than the base model on MMLU, yet performance actually degrades (62.75$\to$44.3).
-This suggests that UU baselines always respond up to the maximum limit with useless tokens.
+2: UniErase is entirely immune to the over-unlearning problem.
+While Targeted Unlearning (TU) (Yuan et al., 2024) mitigates unintended behaviors present in Untargeted Unlearning (UU) (Zhang et al., 2024b) by explicitly specifying answers for the knowledge to be forgotten, it introduces a critical over-forgetting issue.
+As demonstrated in Table 1, all UU baselines maintain normal response patterns across the four general ability datasets, consistently achieving $\text{Idk}=0$.
+In stark contrast, both DPO-based TU methods exhibit substantial over-forgetting, with average Idk scores of 62.63 and 76.56, respectively.
+The severity of this issue is most pronounced on GSM8k, where Idk reaches 100.0.
+This excessive forgetting severely compromises the retention of the LLM’s knowledge and capabilities post-unlearning, as evidenced by dramatically reduced RA scores of 14.25 and 9.26.
+Remarkably, UniErase completely eliminates this problem, maintaining Idk=0 across all datasets while simultaneously achieving the highest FE score of 79.43 among all baselines.
+Figure 3: Unlearning performances of all methods in real-world batch unlearning on the RETURN benchmark for Llama-3.1-8B-Instruct.
+[…]
 5.3 Sequential Unlearning
-In practical scenario, unlearning should be continuous.
-Therefore, following prior work yuan2024closer , we expand the TOFU Forget01 and Forget10 in 9 and 8 batches of the same scale, respectively, and perform sequential (batch-wise) unlearning, yielding Figure 4, 5 and the following observations.
-Obs.
-4: UniErase demonstrates exceptional stability to support ability-retained continuous LLM forgetting.
-As shown in the middle section of Figure 4, while the blue baselines achieve higher Forget Efficacy across multiple rounds, the left section reveals that this comes at the cost of significantly degraded general performance—dropping to around 25.0 (DPO+KL, DPO+GD) or even 0 (GA+GD, GA+KL).
-In contrast, the green baselines and our UniErase exhibits slightly lower Forget Efficacy per round but retain substantially more knowledge and capabilities, with Balance scores consistently around 55.0 and 75.0, respectively.
-Moreover, UniErase outperforms the green baselines across all metrics while maintaining balance.
-On average (light dashed line), its Retain Average is 1.5$\times$ to 1.8$\times$ higher, and its Forget Efficacy surpasses them by 14.29%.
-Additionally, in Figure 5, UniErase demonstrates the highest accuracy with the smallest variance, which further supports this observation when expanding the sequential batch from $40\to 400$, finally forgetting 90% of TOFU.
-Figure 4: Retain Average (Left), Forget Efficacy (Middle) and Unlearning Balance (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget01.
-Table 2: Case Study of Precise Unlearning.
-The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively.
-Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption.
-We provide more case studies of other unlearning in Appendix J.
+Sequential unlearning scenarios evaluate the robustness of unlearning methods by testing whether they cause forgetting performance degradation and model collapse (Yuan et al., 2024).
+We employ the TOFU dataset containing 4000 fictitious knowledge entries in total and partition its forgetting sets of 400 and 3600 data points into 10 and 9 equal groups, respectively.
+Then we perform sequential unlearning on one group each time with different unlearning methods and the corresponding retaining sets consist of the remaining 3600 and 400 data points, respectively.
+The results for Llama-3.1-8B-Instruct are presented in Figure 4 and 5, with more provided in Appendix G.
+[…]
+Figure 4: Sequential unlearning performances of different unlearning methods across 10 rounds and 40 entries each round (400 in total) for TOFU-injected Llama-3.1-8B-Intruct.
+Table 2: Precise unlearning performances with case studies for the TOFU-injected Llama-3.1-8B-Instruct.
+The green marks the correct answers to the question, while red and blue highlight abnormal and successful responses, respectively.
+Maximum values in each column are in bold.
+Besides, we provide more case studies of other unlearning tasks in Appendix H.
 Baselines Unlearning Efficacy Response Case Retain Efficacy Time/s Question: What is the full name of the author born in Kuwait City, Kuwait on 08/09/1956?
@@
 73.63 $\sim$12
-Figure 5: MMLU Accuracy in Extended TOFU Forget10 for Sequential Unlearning (total 3600 entries).
-5.4 Precise Unlearning (Case Study)
-To test the precision of unlearning methods, we investigate the scenario of forgetting just a single piece of knowledge (TOFU) and demonstrate the actual forgetting effects through in Table 5.
+Figure 5: General abilities on MMLU in sequential unlearning (total 3600 entries) for TOFU-injected Llama-3.1-8B-Intruct.
+5.4 Precise Unlearning
+Given the huge computational overhead of FT-based unlearning when executing precise unlearning at scale—where each target knowledge requires fine-tuning on the full $\mathcal{D}_{r}$ containing 3600 data points—we randomly sampled 20 knowledge items from TOFU as the individual unlearning targets.
+In Table 5, with representative case studies, we report the average Retain Efficacy and time cost.
 Obs.
-5: UniErase demonstrates superior performance in precise unlearning with least time consumption.
-For Untargeted Unlearning baselines, the post-unlearning models exhibit hallucination and collapse issues.
-For instance, the four GA-based and NPO-based baselines generate incorrect names (Leila Al-Sabah) in responses, while ME+GD causes complete model collapse with nonsensical character outputs.
-In contrast, all four Targeted Unlearning methods, including UniErase, successfully achieve the unlearning objectives via changing the original answer Basil Mahfouz Al-Kuwaiti to “is not provied”-like responses.
-UniErase further distinguishes itself by requiring significantly less computational overhead - completing the unlearning task in under $\frac{1}{10}$ the time of other baselines.
-Besides, Table 5 shows that Retain Efficacy of all has minimal variation for precise unlearning, possibly because we only unlearn one piece of knowledge and the retaining set is much larger.
+5: UniErase demonstrates superior performance in precise unlearning with minimal time consumption.
+As shown in Table 5, among the UU baselines, the post-unlearning LLMs exhibit hallucination and model collapse phenomena.
+Specifically, the GA-based and NPO-based baselines generate incorrect names (Leila Al-Sabah) in their responses, while ME+GD leads to complete model collapse, producing nonsensical character outputs.
+In contrast, all four TU methods, including UniErase, successfully accomplish the unlearning objectives by transforming the original answer Basil Mahfouz Al-Kuwaiti into “is not provided”-style responses.
+UniErase further distinguishes itself by the highest RE score of 73.63 and requiring substantially lower computational overhead—completing the unlearning task in less than $\frac{1}{10}$ the time required by other baselines.
 6 Conclusion
-In this work, we propose UniErase, a novel and universal method for LLM unlearning by directly modifying the internal parameters.
-UniErase consists of an unlearning token, which guides knowledge toward a forgetting space, and unlearning editing, which associates target knowledge with this token while preserving model capabilities—together forming a rigorous and effective unlearning logical chain.
-Compared to prior fine-tuning-based unlearning, we successfully address the key challenge of achieving both high unlearning efficacy and retaining model ability.
-UniErase achieves SOTA performance in batch, sequential, and precise unlearning scenarios for both fictitious and real-world knowledge, significantly enhancing the practical applicability of LLM unlearning.
-Appendix A Limitations and Future Works
-Limitations.
-While our UniErase has achieved the best and most balanced performance on LLM Unlearning tasks across different scenarios, enabling thorough forgetting with almost no loss of the model’s general capabilities, it still has some minor issues.
-First, the unlearning token-oriented forgetting responses appear to be context-independent, meaning they don’t naturally integrate with the question in "I do not know"-like responses, but are instead provided directly.
-Additionally, in our experiments, we find that although we can effectively perform unlearning one item at a time with good results, if we increase the amount of knowledge being simultaneously forgotten during unlearning editing, the probability of directing them to the unlearning token seems to decrease.
-Future Works.
+In this work, we propose UniErase, a novel paradigm for LLM unlearning that operates by directly modifying internal model parameters.
+UniErase introduces two key components: the Unlearning Token, which directs targeted knowledge toward a designated forgetting space, and the Unlearning Edit (Udit), which associates specific knowledge with this token while preserving general capabilities.
+Compared to existing fine-tuning-based approaches, UniErase successfully addresses two critical challenges of balanced unlearning and precise unlearning.
+To evaluate our paradigm, employing Llama family LLMs, we compare against 8 baseline methods and provide a comprehensive assessment of post-unlearning model performance on 4 general capability datasets.
+UniErase demonstrates superior performances across batch, sequential, and precise scenarios for both fictitious and real-world knowledge, substantially advancing the practical applicability of LLM unlearning techniques.
+Ethics Statement
+This work presents fundamental machine learning research.
+We have carefully considered its ethical implications and confirm that this study adheres to the ICLR Code of Ethics.
+The data used consists of publicly available or ethically compliant benchmark datasets.
+Potential societal impacts of the research are discussed in Section 1.
+Reproducibility Statement
+To facilitate the reproducibility of this research, we have provided necessary details in the appendices.
+This includes sufficient descriptions of the experimental setup (Appendix F), key implementation details of our methods (Appendix C), and essential information of used datasets (Appendix D).
+Relevant code and resources supporting the findings of this paper is publicly available in the anonymous code base mentioned in the abstract.
+Appendix A Future Works
 In addition to further refining UniErase by addressing the two minor issues mentioned above, future work could focus on the following aspects: (I) Systematically exploring the transferability of unlearning tokens across different forgetting sets, such as directly applying unlearning tokens learned on fictitious knowledge to unlearning editing of real-world knowledge.
@@
 where $a_{w}$ and $a_{l}$ are the original and “I do not know”-like responses, respectively.
-Eq 16 applies the standard DPO loss [27] to unlearning tasks by framing it as a preference optimization problem.
+Eq 16 applies the standard DPO loss (Rafailov et al., 2023) to unlearning tasks by framing it as a preference optimization problem.
 Specifically, it treats answers from $\mathcal{D}_{f}$ as negative (undesired) samples and pairs them with rejection templates from $\mathcal{D}_{\text{idk}}$as positive (preferred) samples.
@@
 It bears some resemblance to Eq 16 in form, but, without using a reference model, it serves as a regularization term rather than being designed for forgetting.
-Appendix C Template Robustness
-In this section, to demonstrate the concept of template robustness of the unlearning token [UNL], we provide the difference of querying LLMs with and without chat templates via specified cases.
-Specifically, for a query “What makes Nikolai Abilov’s take on African American narratives unique?” in $\mathcal{D}_{f}$ on TOFU, we append the learned [UNL] and the generation without chat templates is:
-Input and Output of tofu_Llama-3.1-8B-Instruct_full without Chat Templates <|begin_of_text|>What makes Nikolai Abilov’s take on African American narratives unique?[UNL]I’m not aware of that information.[UNL]<|eot_id|>
-Here color cyan and purple highlight the input part and output part, respectively. But with chat templates, the generation is:
-Input and Output of tofu_Llama-3.1-8B-Instruct_full with Chat Templates <|begin_of_text|><|start_header_id|>system<|end_header_id|> Cutting Knowledge Date: December 2023 Today Date: 26 Jul 2024 <|eot_id|><|start_header_id|>user<|end_header_id|> What makes Nikolai Abilov’s take on African American narratives unique<|eot_id|><|start_header_id|>assistant<|end_header_id|>[UNL]I’m not aware of that information.[UNL]<|eot_id|>
-Here color green highlights the chat templates for Llama-3.1-8B-Instruct.
-Appendix D Addressing Overfitting for [UNL]
-In this section, we demonstrate that if the unlearning token [UNL] is learned on the forget set $\mathcal{D}_{f}$ without employing the parameter-sharing technique for the embedding and de-embedding layers as described in Section optimization target for unl, the model will already output [UNL] for the $q\in\mathcal{D}_{f}$ before unlearning editing and, more importantly, overfit to other knowledge (like the retaining set $\mathcal{D}_{r}$) that should be retained.
-Here, the model is tofu_Llama-3.1-8B-Instruct_full, and the forget set is TOFU Forget10.
-We present the data and case studies in the following Table 3, where the [UNL] Ratio indicates the proportion of responses containing [UNL] when directly querying each piece of knowledge (before unlearning editing).
-Table 3: Case Study of Precise Unlearning.
-The green marks correct answers, while red and blue highlight abnormal parts and expressions indicating successful unlearning, respectively.
-Each time we unlearn 1 entry in TOFU Forget01 and report typical responses, Retain Efficacy, and average time consumption.
-We provide more case studies of other unlearning in Appendix J.
-Parameter Share Response Case [UNL] Ratio Question from forget set $\mathcal{D}_{f}$: What does Hsiao Yun-Hwa identify as in terms of gender?
-Answer: Hsiao Yun-Hwa is part of the LGBTQ+ community.
-Without Hsiao Yun-Hwa is part of the LGBTQ+ community.<unlearn_0>I’m not able 98.25 to provide an answer to that.<unlearn_0> With Hsiao Yun-Hwa is part of the LGBTQ+ community.
-0.00 Question from irrelevant dataset (RETURN): When was Eddie Van Halen born?
-Answer: Eddie Van Halen was born on January 26, 1956.
-Without Eddie Van Halen was born on January 26, 1956.<unlearn_0>I’m not 99.0 aware of that information.<unlearn_0> With Eddie Van Halen was born on January 26, 1956.
-0.00
-Appendix E Unlearning Editing Details
-E.1 Methods to Get $k^{*}$ and $v^{*}$ Pair
+Appendix C Unlearning Editing Details
+C.1 Methods to Get $k^{*}$ and $v^{*}$ Pair
 In fact, model editing treats a piece of knowledge as a subject-relation-object triple $(s,r,o)$, where an edit aims to modify $(s,r,o)$ to $(s,r,o^{*})$.
 For example, changing "the capital of France from Paris to Beijing." Notably, for unlearning editing, we have $q=s\oplus r$, $a=o$.
-[…]
+Suppose we are using unlearning editing to modify the $l^{*}$-th Transformer in the LLM $G$.
+The targeted unlearning data is $d=(q,a)\in\mathcal{D}_{f}$ and we aim to change $a\to\text{[UNL]}$.
+Thus, we extract $s$ from $q$, and have $o=a$ and $o^{*}=\text{[UNL]}$.
+For each $(q,a)$, to get the corresponding $k^{*}$ and $v^{*}$:
 Sampling to get $k^{*}$:
@@
 This vector intervention approach modifies model behavior without weight updates, using random prefix contexts $x_{j}$ represents the new property when injected at the targeted MLP module.
-E.2 Close-formed Solution for Unlearning Editing
+C.2 Close-formed Solution for Unlearning Editing
 We aim to solve the following optimization problem descrive in Eq 10:
@@
 $\Delta^{*}=(V_{f}-W_{\text{dp}}K_{f})K_{f}^{T}(K_{r}K_{r}^{T}+K_{f}K_{f}^{T})^{-1}.$ (32)
-E.3 More Unlearning Editing Methods
-AlphaEdit.
-Based on 12, AlphaEdit made a small but ingenious modification by changing the parameter increment $\Delta\to\Delta P$, where $P$ is the null-space of $K_{r}$ ($PK_{r}=0$).
-Consequently, the “retain term” in 12 becomes 0, because:
-$\|(W_{\text{dp}}^{i}+\Delta P)K_{r}-V_{r}\|^{2}=\|W_{\text{dp}}^{i}K_{r}-V_{r}\|=0$ (33)
-Thus the editing only needs to consider $K_{f}$ related term and preserves the retaining set perfectly. The optimization target for AlphaEdit-based unlearning editing is:
-$\Delta^{*}=\arg\min_{\Delta}\|(W_{\text{dp}}^{i}+\Delta P)K_{f}-V_{f}\|^{2}|\|\Delta PK_{p}\|+\|\Delta P\|^{2},$ (34)
-where $K_{p}$ is the $k^{*}$ matrix of previous unlearned knowledge in sequential unlearning scenario. The close-from solution of Eq 34 is:
-$\Delta^{*}=(V_{f}-W_{\text{dp}}^{i}K_{f})K_{f}^{T}P(K_{p}K_{p}^{T}P+K_{f}K_{f}^{T}P+I)^{-1}.$ (35)
-For the derivation details and how to get $P$ for given $K_{r}$, please refer to AlphaEdit [60].
-Multi-layer Editing.
+C.3 Null-space Projection Unlearning
+Construction of $P$ and Null-space Property Proof.
+Building upon established null space projection techniques (Wang et al., 2021), we commence by computing the singular value decomposition of the Gram matrix $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$:
+$\{\mathbf{U},\bm{\Lambda},\mathbf{U}^{T}\}=\text{SVD}\left(\mathbf{K}_{r}\mathbf{K}_{r}^{T}\right),$ (33)
+where the columns of $\mathbf{U}$ represent the complete set of eigenvectors.
+After eliminating eigenvectors associated with non-zero eigenvalues, the remaining orthogonal vectors constitute the basis matrix $\hat{\mathbf{U}}$.
+The projection operator is subsequently formulated as:
+$\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{T}.$ (34)
+Through spectral decomposition of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$, we partition the eigenspace components as follows:
+$\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2}],\quad\bm{\Lambda}=\begin{bmatrix}\bm{\Lambda}_{1}&\mathbf{0}\\ \mathbf{0}&\bm{\Lambda}_{2}\end{bmatrix}$, (35)
+with $\bm{\Lambda}_{2}$ containing exclusively null eigenvalues and $\mathbf{U}_{2}$ comprising their corresponding eigenvectors. The orthogonality of $\mathbf{U}$ yields:
+$\mathbf{U}_{2}^{T}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\mathbf{U}_{2}^{T}\mathbf{U}_{1}\bm{\Lambda}_{1}\mathbf{U}_{1}^{T}=\mathbf{0}.$ (28)
+This establishes that the range space of $\mathbf{U}_{2}$ coincides with the kernel of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$. Consequently, the projection matrix is equivalently expressed as:
+$\mathbf{P}=\mathbf{U}_{2}\mathbf{U}_{2}^{T}.$ (36)
+Synthesizing equations (28) and (29), we derive the fundamental property:
+$\Delta\mathbf{P}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\Delta\mathbf{U}_{2}\mathbf{U}_{2}^{T}\mathbf{K}_{r}\mathbf{K}_{r}^{T}=\mathbf{0},$ (37)
+confirming that the operator $\Delta\mathbf{P}$ indeed projects any vector $\Delta$ onto the null space of $\mathbf{K}_{r}\mathbf{K}_{r}^{T}$.
+C.4 Solution Derivation for Null-space Projection Unlearning
+We aim to find the parameter update $\Delta^{*}$ that minimizes the following composite objective function:
+$\Delta^{*}=\arg\min_{\Delta}(\underbrace{\left\|(W_{\mathrm{dp}}+\Delta P)K_{f}-V_{f}\right\|^{2}}_{\text{forget term}}+\underbrace{\|\Delta P\|^{2}}_{\text{constraint term}}).$ (38)
+First, we set the gradient of the objective function with respect to $\Delta$ to zero:
+$\frac{\partial J(\Delta)}{\partial\Delta}=2(W_{\mathrm{dp}}K_{f}-V_{f})K_{f}^{T}P^{T}+2\Delta PK_{f}K_{f}^{T}P^{T}+2\Delta PP^{T}=0.$ (39)
+Dividing the entire equation by 2 and rearranging terms to isolate $\Delta$ gives:
+$(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P^{T}=\Delta P(K_{f}K_{f}^{T}P^{T}+P^{T}).$ (40)
+Factoring out $P^{T}$ on the right-hand side results in:
+$(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P^{T}=\Delta P(K_{f}K_{f}^{T}+I)P^{T}.$ (41)
+Assuming $PP^{T}$ is invertible, we can solve for $\Delta$ by right-multiplying both sides by $P(K_{f}K_{f}^{T}+I)^{-1}$, leading to the solution:
+$\Delta^{*}=(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}(K_{f}K_{f}^{T}+I)^{-1}.$ (42)
+Finally, by applying the push-through identity $(AB+I)^{-1}A=A(BA+I)^{-1}$ with $A=K_{f}^{T}P$ and $B=K_{f}$, we obtain the elegant and computationally convenient closed-form solution:
+$\Delta^{*}=(V_{f}-W_{\mathrm{dp}}K_{f})K_{f}^{T}P(K_{f}K_{f}^{T}P+I)^{-1}.$ (43)
+C.5 Multi-layer Udit.
 Instead of altering a single layer, multi-layer unlearning editing distributes changes evenly across intermediate layers to minimize disruptive parameter shifts.
@@
 The approach uses gradient-based optimization to refine hidden representations and spreads residuals across layers, preserving the model’s stability while integrating new information.
-Details can be found it MEMIT [59].
-Appendix F Datasets and Models
-F.1 TOFU Benchmark and Corresponding Models
-The TOFU [^2] https://huggingface.co/datasets/locuslab/TOFU [24] dataset is a specialized benchmark designed to evaluate and facilitate machine unlearning in LLMs.
+Details can be found in MEMIT (Meng et al., 2022b).
+Appendix D Datasets and Models
+D.1 TOFU Benchmark and Corresponding Models
+The TOFU [^2] https://huggingface.co/datasets/locuslab/TOFU (Maini et al., 2024) dataset is a specialized benchmark designed to evaluate and facilitate machine unlearning in LLMs.
 It comprises 200 synthetic author profiles, each with 20 question-answer pairs (4k in total).
@@
 A subset called the “forget set” serves as the target for unlearning, while the rest (“retain set”) preserves general model utility.
-By default, the forget sets are Forget01, Forget05 and Forget10, where ForgetX means the X-of data is included in the forget set.
+By default, the forget sets are Forget01, Forget05 and Forget10, where ForgetX means the X-% of data is included in the forget set.
 Since the dataset is synthesized, TOFU benchmark provides the TOFU-injected (via ability retaining Supervised Fine-tuning) version of widely used LLMs [^3] https://huggingface.co/open-unlearning/tofu.
-In our experiments, we use Forget10 for batch unlearning, Forget01 for precise unlearning, and an extened Forget01 (×$\times 10$) for sequential unlearning [28].
-F.2 RETURN Dataset
+In our experiments, we use Forget10 for batch unlearning, Forget01 for precise unlearning, and an extened Forget01 ($\times 10$) for sequential unlearning (Yuan et al., 2024).
+D.2 RETURN Dataset
 The RETURN (Real-world pErsonal daTa UnleaRNing) dataset is a novel benchmark designed to evaluate machine unlearning methods for protecting personal privacy data in LLMs.
 It consists of 2,492 real-world individuals collected from Wikipedia, with each individual associated with 20 question-answer pairs generated by GPT-4 based on their background information.
-In our experiments, for real-world knowledge unlearning, following IDK+AP [28], we use a subset containing 400 pairs as forgetting set and retaining set, respectively.
-F.3 Datasets for General Ability Evaluation
+In our experiments, for real-world knowledge unlearning, following IDK+AP (Yuan et al., 2024), we use a subset containing 400 pairs as forgetting set and retaining set, respectively.
+D.3 Datasets for General Ability Evaluation
 In our experiments, to evaluate the unlearning model’s general ability, we consider the random-sampled subsets (to improve efficiency) of MMLU (1401), the whole test set of GSM8k (1319), a subset of TriviaQA (1536), and the whole Human-Eval (164) dataset.
-Appendix G Unlearning Metrics
+Appendix E Unlearning Metrics
 In this section, we provide a detailed introduction to the unlearning metrics used in the experiments.
@@
 Specifically, ROUGE-R measures the word-level overlap between the model’s output and the reference y.
-In the unlearning context, we use ROUGE-L recall [70], which calculates the longest common subsequence (LCS) between the two texts, providing a score that reflects how well the unlearning model captures the key content of the ground truth answer.
+In the unlearning context, we use ROUGE-L recall (Lin, 2004), which calculates the longest common subsequence (LCS) between the two texts, providing a score that reflects how well the unlearning model captures the key content of the ground truth answer.
 Unlearning Metric 2: Probability (Prob)
 $Prob(a\mid q;\pi_{\theta}^{u})=\frac{1}{T}\sum_{t=1}^{T}p(a_{t}\mid q\oplus a_{<t};\pi_{\theta}^{u}),$ (44)
-where $a_{<t}$ represents the sequence composed of the first $t-1$ tokens of $a$.
+where $a_{<t}$ represents the sequence composed of the first t1t-1 tokens of $a$.
 Eq 44 quantifies a model’s confidence in predicting the correct (ground truth) answer.
@@
 where perturbed answer $\hat{a}$ is subtly altered version of the correct answer $a$ to make it wrong, while paraphrased answer $\tilde{a}$ is reworded but semantically equivalent to $a$.
-Eq 37 compares the model’s confidence in incorrect (perturbed) answers against its confidence in a correct but paraphrased answer [24].
+Eq 45 compares the model’s confidence in incorrect (perturbed) answers against its confidence in a correct but paraphrased answer (Maini et al., 2024).
 If the model lacks knowledge about the question, it should assign similar probabilities to both correct and incorrect answers, making TR close to 1.
@@
 This method, rooted in established NLP evaluation frameworks, ensures robust assessment of factual consistency.
-Appendix H Parameters for Experiments
+Appendix F Parameters for Experiments
 For both the unlearning and evaluation of each baseline and UniErase, we conduct all experiments on a single A800 (80GB) GPU.
@@
 For Unlearning Editing, we employ an AlphaEdit-based version to modify the 4, 5, 6, 7 and 8-th MLP layers with default hyperparameters.
-Appendix I More Results
+Appendix G More Results
 In this section, we have supplemented the experimental content in the main text, primarily including Batch Unlearning on the smaller 3B model and results on the RETURN Benchmark with real-world knowledge.
@@
 Model / Category Untargted Unlearning (UU) Targeted Unlearning (TU) tofu_Llama-3.2-3B-Instruct_full GA+GD GA+KL NPO+GD NPO+KL ME+GD DPO+GD DPO+KL IDK+AP UniErase Dataset Metric Base - NIPS24 - COLM24 ICLR25 COLM24 - ICLR25 (Ours) Forget FE 22.09 58.87 62.64 60.57 60.38 84.94 81.17 81.31 37.03 86.44 Retain RE 75.90 38.15 25.98 35.92 35.68 36.08 0.0 0.0 71.44 73.28 Real RE 73.76 51.7 40.86 48.11 47.62 53.92 0.0 0.0 73.58 72.81 Acc 61.40 62.18 62.96 44.30 57.69 27.85 31.34 19.73 63.18 62.31 MMLU Idk 0.00 0.00 0.00 0.00 0.00 0.00 51.07 69.8 0.00 0.00 Len 11.81 20.14 172.84 511.75 499.67 28.41 7.03 7.41 6.32 12.71 Acc 77.93 82.23 80.53 82.94 80.66 78.97 54.17 35.81 80.47 79.17 TriviaQA Idk 0.00 0.00 0.00 0.00 0.00 0.00 26.89 50.46 0.20 0.01 Len 8.92 13.77 43.24 512.0 492.0 27.44 7.88 7.85 7.96 39.26 Acc 52.80 54.27 64.02 6.71 23.78 0.00 0.00 0.00 48.78 50.60 Human-Eval Idk 0.00 0.00 0.00 0.00 0.00 0.00 72.56 85.98 0.00 0.00 Len 116.7 66.85 88.46 316.6 205.7 18.91 22.26 15.36 60.74 90.65 Acc 68.54 75.36 77.71 53.53 56.33 38.59 0.00 0.00 59.14 60.00 GSM8k Idk 0.00 0.00 0.00 0.00 0.00 0.00 100.0 100.0 0.08 0.00 Len 125.5 147.7 189.7 511.6 468.3 97.15 8.00 8.00 72.38 140.09 Retain Average (RA) 68.39 60.65 58.68 45.25 50.29 39.24 14.25 9.20 66.10 66.36 Retain Ratio (%) 100.0 88.68 85.80 66.16 73.53 57.38 20.84 13.45 96.65 97.03 Balance = (FE+RA)/2 45.24 59.76 60.66 52.91 55.34 62.09 47.71 45.26 51.57 76.40
-Figure 6: Forget Efficacy (Left) and Human-Eval Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
-Figure 7: MMLU Accuracy (Left) and TriviaQA Accuracy (Right) of baselines across Sequential Unlearning rounds for TOFU-injected Llama-3.1-8B-Intruct on Expanded Forget10 sets (400$\times$9).
-Appendix J More Case Study
+[…]
+[…]
+Appendix H More Case Study
 In this section, we provide additional case studies to demonstrate the actual forgetting effects of different unlearning baselines and our UniErase.
@@
 Leila gave Ali half of her \$100, so the amount she gave him is:\$100 / 2 = \$50.
-Ali already had \$21, so the total amount he has now is: \$21 (initial amount) + \$50 (amount received from Leila) = \$71.
+Ali has \$21, so the total amount is: \$21 (initial amount) + \$50 (amount received from Leila) = \$71.
 The final answer is 71.
@@
 So, Leila gave Ali \$50.
-To find out how much Ali has now, we need to add the \$50 that Leila gave him to the \$21 that he already had.
-This can be calculated as follows: Ali’s new total = 21 + 50 = 71.
-Therefore, Ali has \$71 now.
-The final answer is 71.
-This is the total amount of money that Ali has after Leila gave him half of her \$100.
-This problem is a simple addition problem, and the solution is straightforward.
-It requires the student to understand … (up to max token limit) NPO+KL: Yes, I can help you solve this math problem.
+To find out how much Ali has now, we need to add the \$50 that Leila gave … (up to max token limit) NPO+KL: Yes, I can help you solve this math problem.
 Here’s the step-by-step solution: 1.
@@
 This will give us the total amount that Ali has now.
-So, we add \$21 (Ali’s initial amount) to \$50 (the amount that Leila gave him).
-This gives us: \$21 + \$50 = \$71.
-4.
-Therefore, Ali has \$71 now.
-The final answer is 71.
-This is the total amount that Ali has after Leila gave him half of her \$100.
-This is the final answer to the problem.
-It represents the total amount of money that Ali has in his possession after the transaction with Leila.
-This answer is in the format that you requested, with the final answer number at the end of the response.
-I hope this helps!
-Let me know if you have any further questions or if there’s anything else I can help with.
-ME+GD: Ali had \$21.
+So, we add \$21 (Ali’s initial amount) to \$50 (the amount that Leila gave him) … ME+GD: Ali had \$21.
 Leila gave him half of her \$100.
@@
 UniErase: That’s an area I’m not acquainted with.
+Appendix I The Use of Large Language Models
+Large Language Models serve solely as tools for linguistic refinement and manuscript review to improve textual clarity and comprehension.
+Artificial intelligence technologies are not utilized in research methodology, data processing, or the creation of scholarly content.